A Complete Bibliography of the Journal of Number Theory (1970–1979)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

16 July 2020 Version 1.00

Title word cross-reference

(5, 5) [318]. (k, r) [240]. (p − 1)2^{−n} [506]. −1 [87, 174]. −pq [139]. 1 [11, 159, 175, 257, 319]. 163^{2} [362]. 2 [95, 116, 128, 204, 209, 211, 343]. 24 [175]. 3 [106, 152, 260, 261, 335, 504]. 4(2p + 1) [328]. 4 mod 8 [354]. 8 [106]. A^{3} + B^{3} + C^{3} + D^{3} = 0 [316]. A_{5} [526]. α_{k} [256]. B_{2} [124]. mod1 [149]. modk [19, 76, 84]. modm [270]. modp [330]. by^{2} + p^{n} = x^{3} [305]. χ [7, 259]. Δ [291]. η(τ) [30]. F [172, 251]. f(n) = ∑_{i=0}(−1)^{i}(i^{n−2i}) [258]. f(X) [407]. f(X^{p^{2r}} − aX^{p^{r}} − bX) [448, 488]. f(X^{p^{r}} − aX) [407]. F_{p^{n}} [407]. F_{p^{n}}(X) [407]. F_{p^{s}} [448, 488]. F_{Π} [241]. G [333]. Γ [371]. k [117, 126, 144, 265, 323, 346, 400, 470]. k(x) [467]. L [24, 44, 71, 96, 245, 349, 393, 509, 514]. L(1, χ) [7]. L(s, χ) > 0 [259]. L_{p}(1, X) [399]. m [65, 373, 437]. Q[G] [177]. GF(2^{n}) [317]. GF(3^{n}) [317]. GF(q) [126]. PSL(2, 7) [491]. n! [429]. n [138, 183, 274, 496, 516]. p [53, 278, 295, 307, 349, 390, 446, 469, 520]. Π [241, 386]. pq [292]. q [72, 274].

Q(−1)^{12} [348]. Q(−3)^{12} [348]. Q(i(11^{1/2})) [456]. r [346]. s = 12 [245]. s > 0 [259]. S_{3} [460]. ∑(x^{2}/yz) [447]. ∑_{k=1}^{3} |2^{π^{k}} − α_{k}| [256]. tan x [184]. τ(n) [382]. x^{2} + 11 = 3^{n} [289]. x^{2} + x + A [281]. x^{2} − 2py^{2} [443]. Ξ [99]. y^{2} + k = x^{3} [111]. y^{2} − k = x^{3} [34]. Z_{3} [372]. Z_{i} [505]. Z_{l} [273]. ζ(s) [493]. ζ_{k}(s) = ζ_{k′}(s) [419].

-adic [53, 209, 278, 295, 307, 349, 390, 469, 520]. -ary [65]. -class [335]. -Closed [333]. -difference [72, 274]. -dimensional [106, 504, 516]. -expansions [172, 251]. -extensions [273, 371, 505]. -function [71, 99]. -functions [24, 96, 349, 393]. -Funktionen [96]. -funtions [514]. -ic [117]. -integers [240]. -invariants [372]. -Monotone [291]. -part [343]. -potency [95]. -Potenzgrad [95]. -prime [346]. -sequences [124]. -series [245, 509]. -ten [437]. -th [44, 126, 265, 437, 470]. -th-power [400]. -tuples [373].

1972 [120, 288]. 1976 [381].

70803 [194].

abelian [51, 105, 191, 232, 292, 295, 318, 383, 388, 487, 497]. abelschen [487]. abelscher [497]. Ableitung [437]. absolute [485]. absolutely [383]. according [531]. accuracy [473]. addition [104, 292, 388]. additive [45, 46, 70, 247, 266]. adic [53, 209, 278, 295, 307, 349, 390, 469, 520]. admissibility [526]. AF [91]. AF-lattices [91]. afforded [232]. algebra [210]. Algebraic [12, 23, 28, 47, 83, 87, 94, 96, 140, 145, 169, 192, 200, 231, 238, 255, 256, 290, 306, 312, 321, 384, 386, 413, 415, 422, 430, 452, 468, 469, 478, 486, 501, 520]. algebraische [430]. algebraischen [94, 306]. Algebraischer [96, 478]. algebras [177, 320]. algorithm [48, 106, 134, 444]. allgemeinen [32]. almost [50, 461]. almost-prime [461]. Ambiguous [232, 398]. analog [127]. analogues [324]. analysis [336]. analytic [228]. angle [455]. Announcement [481]. anomalous [370]. Antisymmetric [125]. Anwendung [30, 113]. any [174]. Anzahl [496]. Anzahlformeln [446]. appear [394]. application [30, 113, 393, 428]. Applications [5, 165, 199, 324, 471, 525, 533]. applied [49]. Approximation [72, 274, 278, 311, 356, 357, 386, 456, 469, 508]. approximations [185, 230]. April [110, 168, 236]. Archimedean [320]. arguments [121]. arising [385, 522]. Arithmetic [15, 18, 74, 95, 98, 154, 181, 271, 391, 428, 467, 473]. arithmetical [75, 85, 143, 197]. arithmetically [479]. Arithmetische [95]. arrangements [291]. Artin [467, 509]. ary [65]. associated [514]. Asymptotic [322, 368, 369, 411, 450, 474]. auftreten [394]. August [27, 82, 130, 186, 254, 308, 350, 414, 463, 510]. Author [55, 101, 156, 225, 285, 325, 378, 439, 482, 532]. Automorphism [211]. automorphisms [218, 248, 255]. automorphs [97, 387]. auxiliary [166]. average [75, 85, 143].

B [478]. B. [478]. Bases [266, 426, 502]. basic [306]. Bateman [369]. Baton [194]. beginnings [301]. behavior [474]. Bellavitis [80]. Bemerkung [363]. Berechnung [64]. Berlowitz [100]. Berndt [100]. Bernoulli [307, 466, 500]. Bernoullischen [500]. Berwick [444]. between [13, 122, 180, 293, 364, 428, 466]. Bhaskaran [533]. bicyclic [113]. bilinear [223]. Binary [37, 42, 63, 78, 123, 139, 188, 204, 206, 208, 221, 387, 438, 453, 468]. biquadratic [113, 311]. biquadratisch [113]. biquadratisch-bizyklische [113]. bizyklische [113]. Board [2, 57, 103, 158, 227, 287, 327, 380, 441, 484]. bodies [95, 408]. body [174]. Bohuslav [381]. bound [283]. bounded [469]. Bounds [18, 265]. Brandler [377]. Buchberger [478]. Buchbergers [478]. Bulygin [203].

calculation [64, 399, 531]. calculations [220]. Caractères [499]. Carlitz [38]. carrés [174]. Cartan [334]. case [34, 402]. cases [195]. Cayley [8, 212, 304]. centers [329]. central [89, 177]. centre [514]. certain [13, 29, 67, 71, 72, 123, 135, 169, 228, 231, 234, 245, 273, 274, 304, 311, 335, 368, 391]. chains [400]. Character [24, 36, 179, 201, 324, 434, 527]. characteristic [116, 211]. characterization [46, 485]. characters [7, 64, 259, 497]. Charakteren [64]. Chevalley [248, 255, 263]. Chowla [31, 77, 81, 182, 190, 297, 314, 412, 512, 517]. circulant [153]. circulants [162]. circular [442]. Class [4, 12, 20, 23, 37, 47, 59, 75, 85, 89, 94, 116, 123, 128, 143, 146, 150, 164, 167, 179, 183, 187, 196, 204, 221, 222, 238, 243, 249, 257, 260, 261, 264, 290, 335, 343, 354, 364, 406, 449, 464, 477, 479, 497, 501]. class-number [204, 222, 257]. classes [19, 76, 84, 235, 272, 398, 415, 522]. classical [49]. Closed [333]. coding [129]. coefficients [83, 422]. cohomology [375]. Combinatorial [181, 332, 337]. come [229]. comme [174]. Common [62]. compact [520]. complete [520]. completion [70]. complex [249, 311, 370, 384, 460]. components [177]. composés [407]. composition [155]. compositum [409, 519]. compound [407]. computer [92]. Computing [169, 442]. concepts [458]. concerning [121, 235, 300, 436]. conditions [67, 410]. conference [194]. Congruence [61, 65, 67, 243, 410]. congruences [263, 307, 351, 364, 500]. congruent [354]. conguruences [480]. conjecture [77, 113, 115, 237, 314, 369, 377, 412, 501]. conjectures [520]. connected [143]. consecutive [73, 142, 265, 418]. constant [33]. constants [401]. Construction [261, 525, 533]. containing [408]. contemporary [513]. continuation [228]. continued [503]. continuous [278]. convergents [50]. convex [242, 358, 408]. corps [174, 407, 448, 488]. Correction [413, 438]. Corresponding [161, 171]. covariances [402]. Covering [33, 114, 344, 348, 351]. Crapo [403]. criterion [187, 389]. critical [58, 403, 509, 514]. Cubic [4, 59, 79, 201, 208, 312, 335, 362, 397, 409, 433, 455, 494, 509, 531]. cubics [317]. curve [52, 229, 242]. curves [69, 112, 192, 355, 370, 514]. cycles [376]. Cyclic [113, 309, 310, 335, 337, 362]. cyclotomic [29, 122, 164, 174, 313, 416, 464, 489]. cyclotomique [174]. cyclotomiques [489]. Cyclotomy [105, 129]. cylinders [237]. cylindres [237].

dans [174]. December [148, 205, 276, 381]. decomposition [429]. Dedekind [58, 253, 375, 404, 420, 437, 507]. Dedekindschen [437]. Dedication [511]. definite [133, 175, 264, 347, 385, 387]. degree [113, 161, 306, 384, 459, 469]. Denominators [341, 420]. d’ensembles [40]. densités [457]. density [10, 411, 506]. d’entiers [457]. derivative [437]. designed [473]. determinant [139]. determinants [216]. Determination [29, 257]. Diagonal [233, 433, 494, 495]. diagrams [51]. Diedergruppe [95]. difference [72, 105, 125, 274, 337]. Differences [330, 425, 476]. differentiability [390]. dihedral [95]. Dimension [175]. dimensional [106, 251, 475, 504, 516]. diophantine [72, 111, 185, 230, 274, 289, 300, 316, 336, 356, 357, 395, 397, 422, 456, 490, 522]. Dirichlet [24, 163, 228, 301, 405]. discriminant [175, 204, 264, 362, 384]. discriminantal [78]. Diskriminante [175]. Distinct [515]. distributed [46]. distribution [19, 50, 60, 61, 76, 84, 149, 159, 234, 294, 302, 331, 368, 374, 465, 504]. distributions [17, 20, 352]. dividing [204]. Divis [381]. divisibility [4, 187, 445]. divisible [152, 183, 260, 261, 530]. Divisor [212, 234, 398]. divisors [19, 62, 74, 76, 78, 84, 151, 281, 360, 361, 418, 496]. d’Olga [25]. domain [33]. domains [253]. dominating [284]. DOTU [363]. DOTU-matrices [363]. DOTU-Matrizen [363]. dritten [306]. Duality [428]. due [190]. d’une [457]. Dwork [49]. dyadic [282]. Dyson [280].

economics [220]. Editorial [2, 57, 103, 158, 227, 287, 327, 380, 441, 484]. Egyptian [134, 341]. eight [167]. eighth [178]. Einheitenberechnung [531]. Einklassige [94]. Einseinheitengruppen [107]. Eisenstein [47, 238, 265]. Elementary [26, 105, 112, 151]. ellipsoids [329]. elliptic [52, 69, 355, 370, 489, 514]. elliptiques [489]. elliptische [52]. Embedding [140, 244, 253]. embeddings [26]. empirical [86]. Empty [159]. endliche [496]. endlichen [293]. endlicher [64]. ensembles [39]. Enumeration [462]. Equal [275]. equation [13, 34, 71, 111, 289, 316, 419]. equations [72, 274, 283, 300, 397, 421, 422, 478, 495]. equivalence [223]. equivalent [479]. Erdos [3, 151, 332, 353, 369, 417, 436]. Erratum [54, 118, 136]. error [9, 240]. est [407]. estimate [382, 450]. estimates [98, 430]. Euclidean [48]. Euler [129, 324]. Exactly [351]. example [403]. exceptionnels [499]. existence [260, 334]. expansion [134]. expansions [172, 251]. Explicit [16, 261]. exploration [92]. exponential [231]. extension [263, 468]. extensions [161, 171, 232, 273, 293, 335, 342, 371, 383, 467, 505]. extremal [239].

factor [464]. Factorisation [407]. Factorization [8, 224, 304, 407]. Factorizations [268]. Factors [63, 79, 155, 303, 402, 428, 438]. families [105]. Farey [92]. February [1, 56, 102, 157, 226, 286, 326, 379, 440, 483]. Fermat [49, 514]. few [303]. field [23, 79, 87, 89, 145, 200, 272, 293, 355, 362, 383, 407, 409, 448, 456, 488, 509, 525, 533]. Fields [4, 11, 12, 26, 28, 29, 32, 47, 59, 69, 107, 113, 123, 128, 140, 147, 152, 164, 167, 177, 183, 198, 218, 233, 238, 243, 249, 257, 260, 261, 282, 283, 290, 306, 312, 313, 333, 335, 343, 348, 354, 371, 372, 384, 386, 393, 398, 401, 406, 415, 416, 449, 460, 464, 468, 479, 485, 495, 497, 501, 529]. fini [407, 448, 488]. Finite [17, 64, 191, 233, 248, 255, 272, 283, 293, 334, 355, 388, 407, 425, 448, 473, 488, 495, 496, 520]. finitely [46]. Finiteness [116, 300, 345]. first [464]. fixed [214]. footnote [431]. form [200, 220, 387, 421, 443, 506]. Formal [345, 455]. Formel [306, 496]. Formen [52, 94, 175, 394]. forms [37, 42, 52, 63, 78, 83, 94, 109, 123, 132, 133, 138, 139, 153, 154, 160, 175, 188, 194, 196–199, 202, 206, 208, 209, 215, 216, 219, 221, 223, 233, 264, 282, 284, 385, 394, 422, 424, 433, 438, 453, 468, 494, 501, 504, 529]. formula [222, 306, 496]. formulas [324, 446]. four [97]. Fourier [149]. fourth [306]. fractions [134, 341, 503]. Frattini [66]. free [191, 402, 502]. French [25, 40, 159, 174, 178, 237, 407, 448, 488, 489]. Fuchs [417]. full [468]. function [58, 65, 69, 71, 99, 121, 123, 151, 181, 190, 231, 240, 290, 382, 398, 406, 437, 529]. functional [71]. functions [24, 46, 72, 75, 85, 96, 143, 149, 234, 247, 258, 270, 274, 278, 347, 349, 391, 393, 452, 474]. Fundamental [11, 45, 170, 373, 376]. Funktionen [96]. funtions [514]. further [74].

G. [35]. Galois [89, 140, 375, 485, 491]. Galoisian [32, 293]. galoissche [32]. Galoisschen [89, 293]. Gauss [222, 518]. Gaussian [36, 41, 331, 435]. genera [206, 219, 424]. general [32, 268, 458]. Generalization [10, 90, 92, 187, 405, 444, 520]. Generalizations [80, 222]. Generalized [97, 131, 467, 500, 507]. generation [520]. genus [4, 383, 409, 519, 525, 533]. German [30, 32, 52, 64, 89, 94–96, 107, 113, 175, 293, 306, 363, 364, 394, 430, 437, 446, 449, 478, 487, 496, 497, 500, 508, 531]. Geschlechter [94, 113]. Ginzburg [332]. given [408]. Gleichungen [478]. global [198, 529]. Godwin [531]. Good [296]. Grades [306]. gradually [229]. graphs [340]. Grössencharaktere [497]. group [52, 53, 95, 135, 170, 177, 198, 210, 221, 293, 295, 313, 318, 320, 342, 406, 491]. group-theoretical [293]. groups [51, 64, 66, 105, 107, 140, 167, 191, 209, 211, 218, 241, 248, 255, 292, 334, 335, 388, 485, 487]. Grundeinheiten [306]. Gruppen [64, 487]. gruppentheoretische [293].

H. [35]. H.-G. [35]. Hadamard [328]. Hall [492]. Hammersley [475]. Hardy [382]. Hasse [49, 112, 370]. having [152, 351]. Hecke [170, 257]. Hermitian [154, 160]. higher [459]. higher-degree [459]. Hilbert [22, 252, 354]. Hilfe [508]. homology [520]. homotopy [515]. huitième [178]. Huxley [365]. hyper [355]. hyper-elliptic [355]. hyperelliptic [112]. hypothesis [257, 259].

ic [117]. idea [49]. ideal [383]. ideals [23, 232, 359]. identical [301]. identities [80]. identity [517]. II [76, 85, 180, 238, 247, 250, 323, 357, 361, 369, 371, 410, 459, 490]. III [15, 84, 150, 191, 309, 337]. imaginär [449]. imaginär-quadratischen [449]. Imaginary [128, 257, 260, 261, 343, 371, 372, 393, 449]. imaginary-square [449]. imbeddings [333]. Including [402]. indefinite [42]. independence [231]. index [55, 101, 156, 225, 285, 325, 378, 439, 482, 532]. indexing [458, 459]. indices [64]. Indizes [64]. inequalities [395, 490]. inequality [373]. infinite [218]. infinitely [260, 261, 530]. Ingham [451]. inhomogeneous [138]. Integer [305, 355, 435, 447, 458, 459]. integers [19, 20, 36, 44, 70, 76, 84, 98, 104, 109, 144, 169, 210, 240, 265, 279, 295, 296, 303, 323, 346, 368, 392, 418, 426, 435, 468, 476, 486, 502, 506, 528]. Integral [99, 121, 123, 132, 143, 213, 224, 232, 247, 453, 504, 522]. integral-valued [247]. interesting [34]. Intersections [40]. intervals [180]. invariant [49, 406]. Invariants [152, 197, 202, 370, 372, 480, 505]. inversion [149]. invertibility [359]. involving [195]. irrational [503]. Irreducibility [252, 448, 488]. irreducible [13, 407]. Irréductibilité [448, 488]. irréductible [407]. irregular [294, 374]. Irregularities [15, 17, 352]. Irregularity [141]. Isomorphism [53]. isotropic [217]. IV [310]. Iwasawa [298, 480].

J [120]. Jacobi [106, 117, 207, 518]. Jacobsthal [518]. January [120]. July [381]. June [119, 176, 246].

kernel [244]. Klassenzahl [89]. Klassenzahlen [364]. Kloosterman [516]. Kongruenzen [364, 500]. Körper [113]. Körpererweiterungen [293]. Kronecker [401, 415]. kubischen [531]. Kummer [6, 187, 480, 500]. Kummersche [500]. Kuroda [288]. Kurve [52]. Kuzmin [172].

L. [120]. Lagrange [127]. lambda [21]. Large [28, 61, 63, 160, 401, 410, 418, 431, 438]. Lattice [146, 229, 242, 329, 358, 408, 427, 458, 459]. lattice-points [458, 459]. lattices [91]. laws [108, 530]. Legendre [117, 207, 269]. lemma [336]. length [106]. lifted [520]. line [58, 509]. Linear [13, 66, 72, 83, 109, 138, 336, 360, 361, 422]. Local [16, 282, 331, 485]. Localizability [244]. loops [337]. lorsque [407]. Lösen [478]. Louisiana [194]. lower [283].

M. [533]. MacLaurin [324]. Mahler [278, 300]. making [195]. many [260, 261]. maps [293]. March [120]. matrices [208, 212, 215, 224, 235, 295, 328, 363, 522]. matrix [224]. Matrizen [363]. Maximal [191, 217, 266, 411]. May [14, 68, 299, 338, 396, 454, 498]. Mean [514]. measures [473]. Menge [394]. Merten [271]. method [21, 28, 195, 203, 451, 478]. metric [45, 356, 357]. Minima [37, 42, 394]. Minimal [266]. minimum [384]. mixed [395, 490]. modified [10]. Modular [170, 340, 385, 504]. modules [468]. moduli [351]. modulo [61, 159, 263, 446, 462, 465]. modulus [7, 319]. Moments [150]. Monotone [291]. Montgomery [165, 521]. Mordell [34, 120]. most [351]. Müller [38]. multidimensional [329]. multipartites [410]. multiples [59]. multiplication [370]. multiplicative [474]. must [229].

nach [531]. Narkiewicz [115]. natural [351, 513]. near [229]. neofields [337]. Néron [52]. Noetherian [66]. Non [259, 320, 520]. Non-Archimedean [320]. non-complete [520]. non-principal [259]. nonbases [266, 411, 502]. nonexistence [13, 459]. nonhomogeneous [37]. nonnegative [109]. nonnormal [161]. nonpartitionings [217]. nonprincipal [7]. nonresidues [73]. nonsolvable [140]. Nonstandard [192]. nontrivial [13]. nontriviality [273]. nonvanishing [245, 434]. norm [11, 16, 214, 284, 293, 421, 468]. normal [40, 95, 171]. Normalkörper [95]. normaux [39, 40]. Note [31, 60, 74, 115, 138, 257, 259, 317, 349, 352, 365, 451, 527]. Notes [489]. November [43, 93, 288, 315, 366, 423, 472, 523]. Nullstellen [437]. Number [4, 5, 12, 21, 23, 28, 32, 45, 47, 48, 58, 87, 89, 107, 113, 116, 128, 131, 140, 145, 200, 204, 222, 238, 239, 242, 257, 260, 261, 266, 272, 281, 283, 290, 306, 312, 333, 335, 339, 343, 344, 354, 367, 384, 386, 402, 408, 415, 418, 428, 433, 442, 446, 449, 464, 494, 496, 497, 501, 530]. Numberfields [321, 413]. numbers [8, 12, 47, 50, 59, 94, 127, 142, 150, 164, 166, 179, 180, 183, 187, 196, 231, 238, 243, 249, 256, 264, 284, 291, 307, 311, 364, 386, 394, 430, 445, 466, 469, 479, 500, 501, 503, 508, 531]. numerous [195].

observations [86]. occurring [351]. October [137, 193, 267]. odd [121, 506]. Olga [25]. one [94, 251, 465]. One-class [94]. one-dimensional [251]. order [75, 85, 143, 240, 292, 328, 360, 361, 402]. orders [169, 221, 359]. Orthogonal [196, 209, 211, 214, 215]. other [514]. overlattices [309, 310].

P. [3, 353]. packing [344]. Pages [1, 14, 27, 43, 56, 68, 82, 93, 102, 110, 119, 130, 137, 148, 157, 168, 176, 186, 193, 205, 226, 236, 246, 254, 267, 276, 286, 299, 308, 315, 326, 338, 350, 366, 379, 396, 414, 423, 440, 454, 463, 472, 483, 498, 510, 523]. papers [365]. part [343]. partial [486]. partition [65, 80]. Partitionings [217]. Partitions [67, 277, 322, 368, 369, 410, 436]. parts [67]. period [106]. periodic [106]. periodicity [336]. permutations [425]. Perron [106]. Pfister [199]. philosophy [513]. Pieces [167]. point [52, 229, 427]. points [146, 192, 242, 329, 355, 358, 408, 458, 459, 504]. Poisson [324]. polyhedra [427]. polynôme [407]. polynômes [407, 448, 488]. Polynomial [70, 270, 387, 407, 458, 459, 465]. Polynomials [62, 155, 268, 278, 312, 385, 407, 435, 448, 458, 459, 488, 491, 524]. Positive [94, 133, 144, 216, 303, 323, 346, 347, 385, 387, 442, 501]. potency [95]. Potenzgrad [95]. power [24, 26, 204, 265, 400, 403, 417, 462, 470]. powers [44, 126, 263, 275, 395, 429, 490, 528]. Preassigned [36]. Prime [7, 21, 60, 61, 63, 107, 113, 131, 144, 161, 263, 302, 303, 323, 339, 346, 360, 361, 367, 402, 418, 428, 429, 438, 461, 462, 464, 477]. Primes [44, 131, 188, 294, 331, 370, 374]. primitive [60, 86, 302]. Primzahlgrad [113]. Principal [74, 79, 259, 383]. probability [20, 144, 323, 346]. problem [3, 31, 35, 81, 140, 178, 182, 239, 244, 329, 332, 348, 353, 403, 432, 436, 522, 524, 530]. problème [178]. problems [356, 357]. product [138, 224]. products [129, 279]. progression [18, 98]. progressions [15, 74, 271, 428, 473]. Projectors [241]. Proof [71, 77, 112, 151, 260, 262, 278, 314, 339, 382, 434]. proofs [195, 307]. properties [65, 162, 189, 243, 391]. property [61, 184, 277, 445]. Prüfer [173]. Punktgruppe [52]. Pure [4, 59, 306, 409]. pythagoräischer [508]. Pythagorean [508].

Quadratic [11, 18, 37, 42, 52, 73, 78, 91, 94, 107, 108, 123, 128, 132, 133, 139, 147, 152, 153, 167, 175, 177, 183, 188, 189, 194, 196– 198, 201, 202, 204, 206, 208, 209, 215, 216, 219– 221, 243, 249, 257, 260, 261, 269, 333, 343, 345, 364, 371, 372, 385, 387, 393, 394, 406, 424, 425, 443, 453, 458, 468, 501, 529]. quadratische [175]. quadratischen [52, 107, 449]. quadratischer [94, 364, 394]. quantitative [149]. Quasi [219, 424]. Quasi-genera [219, 424]. quaternary [264]. quaternion [210]. Quaternionic [519]. quaternions [213]. quelconque [174]. question [25]. questions [506]. quick [382]. quotients [486].

R [492]. R. [492]. Raikov [5]. Ramachandra [365]. Ramanujan [9, 382]. Random [250, 367]. Rational [4, 52, 295, 347, 409, 508, 515]. Rationale [508]. rationalen [52]. rationals [453]. readable [195]. Real [7, 11, 50, 94, 147, 177, 183, 189, 259, 291, 372, 430, 501, 531]. reciprocity [108, 269]. reconsideration [207]. recurrences [360, 361]. reduction [133, 140]. reelle [430]. reflection [32]. region [170]. regular [91]. Regulatorabschätzungen [430]. Regulatory [430]. reinen [306]. related [22, 154, 231, 289, 389, 450, 506]. relation [466]. Relationen [293]. relations [13, 94, 113, 122, 293, 322]. relativ [89, 113]. relativ-galoisschen [89]. relativ-zyklischer [113]. relative [15, 79, 89, 113, 293, 497]. relative-cyclic [113]. relatively [144, 323, 346, 477]. Relativklassenzahl [497]. Relativnormabbildungen [293]. relevance [513]. Remainder [98, 107]. remark [363]. Remarks [79, 121, 172, 521]. Remarques [489]. répartition [159, 457]. Report [194]. Representation [87, 99, 174, 284, 453, 493]. Representations [78, 109, 126, 132, 232, 468]. representatives [515]. represented [188]. Residuacity [189]. residue [7, 16, 19, 76, 84, 161, 171, 400]. residues [18, 117, 265, 330, 403, 425, 470]. Restklassengruppen [107]. result [166, 235, 492]. results [22, 344, 368, 369, 459]. réunion [39]. Riemann [99, 121]. ring [197, 212, 345, 449, 455, 524]. Ringklassenkörper [449]. Rings [53, 135, 173, 213, 253, 304, 375]. roots [13, 60, 86, 302, 416]. Rota [403]. Roth [452]. Rouge [194]. Round [198, 282].

S. [31, 77, 190, 297, 314]. Salié [88]. Sarvadaman [512]. satisfying [67, 72, 274]. Satz [113]. Schmidt [452]. Schneider [178]. Schnirelmann [10]. Schreier [467]. Schur [313]. schurschen [64]. science [513]. sechsten [306]. Second [360, 361, 402]. Second-order [360, 402]. sector [345]. seine [113]. Selberg [450, 517]. sequence [289]. Sequences [15, 17, 73, 92, 104, 124, 141, 146, 214, 230, 296, 367, 465, 475, 477]. series [26, 163, 228, 245, 301, 417, 509]. set [394]. sets [40, 70, 125, 191, 318, 337, 373, 426, 453]. several [391]. sham [64]. Shimura [81]. short [434]. Siegel [300, 452]. Sierpiński [34]. sieve [28, 165, 431, 450, 461]. sieves [160, 250]. sifted [141]. Sigekatu [288]. sign [41, 404]. significance [181]. similarity [522]. Simple [71, 177, 248, 255, 278, 339, 503]. Simultaneous [274, 495]. singular [520]. six [114]. sixth [275, 306, 384]. size [7, 497]. Skew [153, 215]. Skolem [432]. small [44, 160, 216, 290]. solid [455]. soluble [241]. Solution [178, 316, 432]. solutions [61, 283, 305, 495]. solving [478]. Some [5, 22, 86, 92, 162, 165, 172, 197, 232, 344, 355, 404, 465, 503, 505, 521, 525, 533]. somme [174]. sous [457]. sous-suites [457]. space [91, 114]. spaces [217]. special [337, 355, 386]. spectre [159]. spectrum [159]. spheres [114, 504]. spherical [385]. Spiegelungssatz [32]. Spinor [206]. Splitting [196, 342]. square [264, 272, 402, 417, 449, 502]. square-free [502]. squarefree [98]. squares [69, 87, 97, 142, 174, 180, 203, 262, 347, 442]. star [33]. Stark [90]. State [194]. Steiner [35]. steps [48]. Stirling [445]. strictly [358]. strip [514]. structure [209, 221, 293]. structures [337]. Struktur [293]. stufe [29, 145, 272]. subgroups [66, 334]. subsequences [465]. suite [457]. suites [159, 457]. sum [41, 87, 88, 97, 151, 174, 191, 256, 347, 373, 434]. sum-free [191]. summation [324, 451]. Summen [487]. Sums [24, 44, 69, 142, 179, 180, 203, 275, 303, 318, 387, 404, 416, 420, 462, 476, 487, 507, 516, 518, 527, 528]. supplementary [352]. supports [520]. surfaces [358]. symbol [16]. symbols [117, 207]. symmetric [224]. symplectic [244]. system [397]. Systeme [478]. systems [161, 171, 351, 478].

tails [530]. take [435]. Tate [49, 52]. Tate’schen [52]. Tauberian [163, 389]. Taussky [25]. Teiler [496]. ten [437]. term [9]. terms [402]. Tesselation [392]. th [44, 126, 265, 400, 437, 470]. their [199, 485]. theorem [5, 6, 22, 32, 38, 49, 90, 112, 131, 151, 163, 172, 185, 251, 252, 262, 263, 269, 271, 278, 280, 292, 297, 298, 300, 339, 354, 367, 388, 389, 405, 417, 428, 452, 521]. theorems [45, 100, 140, 383]. theoretical [293]. Theorie [95, 449]. theory [4, 5, 21, 45, 95, 166, 239, 266, 344, 427, 449, 530]. third [306]. three [59, 152, 262, 347]. Thue [452]. tori [96]. total [430]. totally [94, 384, 430, 442, 460, 501, 531]. totalpositiver [94]. totalreellen [94, 531]. trace [200]. trace-form [200]. transcendental [166]. transformation [251]. triangular [142]. tuples [373]. turning [229]. twice [351]. Two [87, 180, 224, 284, 347, 356, 357, 475]. two-dimensional [475]. type [47, 238, 318]. types [152].

unequal [67]. Uniform [465, 504]. Unimodular [209, 215, 282, 529]. unipositive [208]. unit [4, 11, 107, 405, 444, 531]. unitary [210, 218]. unités [489]. Units [122, 135, 147, 169, 189, 201, 306, 319, 321, 376, 413, 442, 460, 471, 489]. unity [416]. University [194]. using [508].

valued [247]. values [36, 62, 72, 121, 234, 274, 435, 447, 514]. variable [274, 278]. variables [284, 347, 391]. variances [402]. varieties [515, 520]. Vaughan [521]. vectors [124, 214]. verallgemeinerten [500]. Verfahren [478]. via [92, 215, 504]. vide [159]. vierten [306]. volume [378, 439, 482, 532]. Voronoi [133].

W. [115]. Waring [524]. Weil [112, 520]. Weyl [389]. which [142, 347, 435, 442]. whose [59, 260, 261]. Witt [197, 253].

Zahlen [394, 500, 508]. Zahlkörper [32, 113, 364, 430, 497]. Zahlkörpern [94, 107, 306, 449, 531]. Zahlkörpers [89]. zentrale [89]. zero [411]. zeroing [437]. Zeros [58, 258, 433, 494, 509]. zeta [58, 121, 437]. zeta-function [58]. Zetafunktion [437]. Ziv [332]. Zur [449, 531]. zwischen [293, 364]. zyklischer [113].

References

Anonymous:1970:PF

[1] Anonymous. Pages 1–123 (February 1970). Journal of Number Theory, 2 (1):??, February 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1970:EB

[2] Anonymous. Editorial Board. Journal of Number Theory, 2(1):i, February 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

     tronic). URL http://www.sciencedirect.com/science/article/pii/
     0022314X70900016.

Katai:1970:PPE

[3] I. Kátai. On a problem of P. Erdős. Journal of Number Theory, 2(1):1–6, February 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

     (electronic). URL http://www.sciencedirect.com/science/article/
     pii/0022314X70900028.

Barrucand:1970:RGC

[4] Pierre Barrucand and Harvey Cohn. A rational genus, class number divisibility, and unit theory for pure cubic fields. Journal of Number Theory, 2(1):7–21, February 1970. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X7090003X.

Elliott:1970:SAT

[5] P. D. T. A. Elliott. Some applications of a theorem of Raikov to number theory. Journal of Number Theory, 2(1):22–55, February 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X70900041.

Chowla:1970:TK

[6] Paromita Chowla. On a theorem of Kummer. Journal of Number Theory, 2(1):56–57, February 1970. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X70900053.

Joshi:1970:SRN

[7] Padmini Tryambak Joshi. The size of L(1, χ) for real nonprincipal residue characters χ with prime modulus. Journal of Number Theory, 2(1):58–73, February 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X70900065.

Pall:1970:FCN

[8] Gordon Pall and Olga Taussky. Factorization of Cayley numbers. Journal of Number Theory, 2(1):74–90, February 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X70900077.

Smith:1970:ETR

[9] R. A. Smith. An error term of Ramanujan. Journal of Number Theory, 2(1):91–96, February 1970. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X70900089.

Freedman:1970:GMS

[10] A. R. Freedman. On the generalization of the modified Schnirelmann density. Journal of Number Theory, 2(1):97–105, February 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X70900090.

Yokoi:1970:FUR

[11] Hideo Yokoi. On the fundamental unit of real quadratic fields with norm 1. Journal of Number Theory, 2(1):106–115, February 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X70900107.

Madan:1970:CNA

[12] Manohar L. Madan. On class numbers of algebraic number fields. Journal of Number Theory, 2(1):116–119, February 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X70900119.

Chowla:1970:NNL

[13] S. Chowla. The nonexistence of nontrivial linear relations between the roots of a certain irreducible equation. Journal of Number Theory, 2 (1):120–123, February 1970. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X70900120.

Anonymous:1970:PM

[14] Anonymous. Pages 125–247 (May 1970). Journal of Number Theory, 2 (2):??, May 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Roth:1970:ISR

[15] K. F. Roth. Irregularities of sequences relative to arithmetic progressions, III. Journal of Number Theory, 2(2):125–142, May 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X70900132.

Geissinger:1970:ELN

[16] L. Geissinger and W. Graves. An explicit local norm residue symbol. Journal of Number Theory, 2(2):143–151, May 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X70900144.

Berlekamp:1970:IDF

[17] E. R. Berlekamp and R. L. Graham. Irregularities in the distributions of finite sequences. Journal of Number Theory, 2(2):152–161, May 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X70900156.

Singh:1970:BQR

[18] Sahib Singh. Bounds of quadratic residues in arithmetic progression. Journal of Number Theory, 2(2):162–167, May 1970. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X70900168.

Hall:1970:DDI

[19] R. R. Hall. On the distribution of divisors of integers in residue classes (modk). Journal of Number Theory, 2(2):168–188, May 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X7090017X.

Golomb:1970:CPD

[20] Solomon W. Golomb. A class of probability distributions on the integers. Journal of Number Theory, 2(2):189–192, May 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X70900181.

Golomb:1970:LMP

[21] Solomon W. Golomb. The lambda method in prime number theory. Journal of Number Theory, 2(2):193–198, May 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X70900193.

Kisilevsky:1970:SRR

[22] H. Kisilevsky. Some results related to Hilbert’s theorem 94. Journal of Number Theory, 2(2):199–206, May 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X7090020X.

Gunji:1970:CIA

[23] Hiroshi Gunji and Donald L. McQuillan. On a class of ideals in an algebraic number field. Journal of Number Theory, 2(2):207–222, May 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X70900211.

Apostol:1970:DFC

[24] Tom M. Apostol. Dirichlet L-functions and character power sums. Journal of Number Theory, 2(2):223–234, May 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X70900223.

Serre:1970:QOT

[25] Jean-Pierre Serre. Sur une question d’Olga Taussky. (French) [On a question of Olga Taussky]. Journal of Number Theory, 2(2):235–236, May 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X70900235.

Robinson:1970:EEF

[26] Abraham Robinson. Elementary embeddings of fields of power series. Journal of Number Theory, 2(2):237–247, May 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X70900247.

Anonymous:1970:PA

[27] Anonymous. Pages 249–378 (August 1970). Journal of Number Theory, 2(3):??, August 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Schaal:1970:LSM

[28] Werner Schaal. On the large sieve method in algebraic number fields. Journal of Number Theory, 2(3):249–270, August 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X70900521.

Chowla:1970:DSC

[29] P. Chowla and S. Chowla. Determination of the Stufe of certain cyclotomic fields. Journal of Number Theory, 2(3):271–272, August 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X70900533.

Wohlfahrt:1970:AGA

[30] Klaus Wohlfahrt. Eine Anwendung von η(τ). (German) [An application of η(τ)]. Journal of Number Theory, 2(3):273–278, August 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X70900545.

Madan:1970:NPC

[31] Manohar L. Madan. Note on a problem of S. Chowla. Journal of Number Theory, 2(3):279–281, August 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X70900557.

Kuroda:1970:ASG

[32] Sige-Nobu Kuroda. Über den allgemeinen Spiegelungssatz für galoissche Zahlkörper. (German) [On the general reflection theorem for Galoisian number fields]. Journal of Number Theory, 2(3):282–297, August 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X70900569.

Hans-Gill:1970:CCS

[33] Rajinder Jeet Hans-Gill. Covering constant of a star domain. Journal of Number Theory, 2(3):298–309, August 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X70900570.

Finkelstein:1970:MEI

[34] Raphael Finkelstein and Hymie London. On Mordell’s equation y^{2} − k = x^{3}: an interesting case of Sierpiński. Journal of Number Theory, 2(3):310– 321, August 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X70900582.

Zassenhaus:1970:PHG

[35] Hans Zassenhaus. On a problem of H.-G. Steiner. Journal of Number Theory, 2(3):322–328, August 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X70900594.

Rabung:1970:PCV

[36] John R. Rabung. Preassigned character values in the Gaussian integers. Journal of Number Theory, 2(3):329–332, August 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X70900600.

Varnavides:1970:NMC

[37] P. L. Varnavides. The nonhomogeneous minima of a class of binary quadratic forms. Journal of Number Theory, 2(3):333–341, August 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X70900612.

Ayoub:1970:TMC

[38] R. Ayoub and S. Chowla. On a theorem of Müller and Carlitz. Journal of Number Theory, 2(3):342–344, August 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X70900624.

France:1970:REN

[39] Michel Mendes France. La réunion des ensembles normaux. Journal of Number Theory, 2(3):345–351, August 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X70900636.

Dress:1970:IEN

[40] François Dress. Intersections d’ensembles normaux. (French) [Intersections of normal sets]. Journal of Number Theory, 2(3):352–362, August 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X70900648.

Waterhouse:1970:SGS

[41] William C. Waterhouse. The sign of the Gaussian sum. Journal of Number Theory, 2(3):363, August 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X7090065X.

Hightower:1970:MIB

[42] Collin J. Hightower. The minima of indefinite binary quadratic forms. Journal of Number Theory, 2(3):364–378, August 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X70900661.

Anonymous:1970:PN

[43] Anonymous. Pages 379–512 (November 1970). Journal of Number Theory, 2(4):??, November 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Prachar:1970:SPP

[44] K. Prachar. On sums of primes and l-th powers of small integers. Journal of Number Theory, 2(4):379–385, November 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X70900405.

Stalley:1970:FMT

[45] Robert D. Stalley. On the fundamental metric theorems of additive number theory. Journal of Number Theory, 2(4):386–392, November 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X70900417.

Ryavec:1970:CFD

[46] C. Ryavec. A characterization of finitely distributed additive functions. Journal of Number Theory, 2(4):393–403, November 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X70900429.

Ishida:1970:CNA

[47] Makoto Ishida. Class numbers of algebraic number fields of Eisenstein type. Journal of Number Theory, 2(4):404–413, November 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X70900430.

Dixon:1970:NSE

[48] John D. Dixon. The number of steps in the Euclidean algorithm. Journal of Number Theory, 2(4):414–422, November 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X70900442.

Chowla:1970:ITD

[49] S. Chowla. An idea of Tate–Dwork for the “Hasse invariant” applied to a classical theorem of Fermat. Journal of Number Theory, 2(4):423–424, November 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X70900454.

Erdos:1970:DCA

[50] P. Erdős. On the distribution of the convergents of almost all real numbers. Journal of Number Theory, 2(4):425–441, November 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X70900466.

Ayoub:1970:DAG

[51] Christine W. Ayoub. On diagrams for abelian groups. Journal of Number Theory, 2(4):442–458, November 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X70900478.

Zimmer:1970:NTQ

[52] Horst Günter Zimmer. Die Néron–Tate’schen quadratischen Formen auf der rationalen Punktgruppe einer elliptische Kurve. (German) [Néron– Tate’s quadratic forms on the rational point group of an elliptic curve]. Journal of Number Theory, 2(4):459–499, November 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X7090048X.

Sehgal:1970:IAG

[53] Sudarshan K. Sehgal. Isomorphism of p-adic group rings. Journal of Number Theory, 2(4):500–508, November 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X70900491.

Anonymous:1970:E

[54] Anonymous. Erratum. Journal of Number Theory, 2(4):509, November 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X70900508.

Anonymous:1970:AI

[55] Anonymous. Author index. Journal of Number Theory, 2(4):511–512, November 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X7090051X.

Anonymous:1971:PF

[56] Anonymous. Pages 1–123 (February 1971). Journal of Number Theory, 3 (1):??, February 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1971:EB

[57] Anonymous. Editorial Board. Journal of Number Theory, 3(1):i, February 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X71900436.

Berndt:1971:NZD

[58] Bruce C. Berndt. The number of zeros of the Dedekind zeta-function on the critical line. Journal of Number Theory, 3(1):1–6, February 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X71900448.

Honda:1971:PCF

[59] Taira Honda. Pure cubic fields whose class numbers are multiples of three. Journal of Number Theory, 3(1):7–12, February 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X7190045X.

Vegh:1971:NDP

[60] Emanuel Vegh. A note on the distribution of the primitive roots of a prime. Journal of Number Theory, 3(1):13–18, February 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X71900461.

Williams:1971:DPS

[61] Kenneth S. Williams. A distribution property of the solutions of a congruence modulo a large prime. Journal of Number Theory, 3(1):19–32, February 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X71900473.

MacCluer:1971:CDV

[62] C. R. MacCluer. Common divisors of values of polynomials. Journal of Number Theory, 3(1):33–34, February 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X71900485.

Greaves:1971:LPF

[63] George Greaves. Large prime factors of binary forms. Journal of Number Theory, 3(1):35–59, February 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X71900497. See correction [438].

Lorenz:1971:BSI

[64] Falko Lorenz. Über die Berechnung der schurschen Indizes von Charakteren endlicher Gruppen. (German) [On the calculation of the sham indices of characters of finite groups]. Journal of Number Theory, 3(1): 60–103, February 1971. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X71900503.

Andrews:1971:CPA

[65] George E. Andrews. Congruence properties of the m-ary partition function. Journal of Number Theory, 3(1):104–110, February 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X71900515.

Sehgal:1971:FSL

[66] Surinder K. Sehgal and Hans Zassenhaus. The Frattini subgroups of linear Noetherian groups. Journal of Number Theory, 3(1):111–114, February 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X71900527.

Hagis:1971:PUP

[67] Peter Hagis. Partitions into unequal parts satisfying certain congruence conditions. Journal of Number Theory, 3(1):115–123, February 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X71900539.

Anonymous:1971:PM

[68] Anonymous. Pages 125–252 (May 1971). Journal of Number Theory, 3 (2):??, May 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Cassels:1971:SSE

[69] J. W. S. Cassels, W. J. Ellison, and A. Pfister. On sums of squares and on elliptic curves over function fields. Journal of Number Theory, 3(2):125– 149, May 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X71900308.

Donagi:1971:ACP

[70] Ran Donagi and Marcel Herzog. On the additive completion of polynomial sets of integers. Journal of Number Theory, 3(2):150–154, May 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X7190031X.

Kitaoka:1971:SPF

[71] Yoshiyuki Kitaoka. A simple proof of the functional equation of a certain L-function. Journal of Number Theory, 3(2):155–158, May 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X71900321.

Osgood:1971:DAV

[72] Charles F. Osgood. On the diophantine approximation of values of functions satisfying certain linear q-difference equations. Journal of Number Theory, 3(2):159–177, May 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X71900333.

Hudson:1971:SCQ

[73] Richard H. Hudson. On sequences of consecutive quadratic nonresidues. Journal of Number Theory, 3(2):178–181, May 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X71900345.

Parry:1971:FNP

[74] Charles J. Parry. A further note on principal divisors in arithmetic progressions. Journal of Number Theory, 3(2):182–183, May 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X71900357.

Berndt:1971:AOCa

[75] Bruce C. Berndt. On the average order of a class of arithmetical functions, I. Journal of Number Theory, 3(2):184–203, May 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X71900369.

Hall:1971:DDIa

[76] R. R. Hall. On the distribution of divisors of integers in residue classes ( mod k), II. Journal of Number Theory, 3(2):204–209, May 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X71900370.

Ree:1971:PCC

[77] Rimhak Ree. Proof of a conjecture of S. Chowla. Journal of Number Theory, 3(2):210–212, May 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X71900382.

Brown:1971:RDD

[78] Ezra Brown. Representations of discriminantal divisors by binary quadratic forms. Journal of Number Theory, 3(2):213–225, May 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X71900394.

Barrucand:1971:RPF

[79] Pierre Barrucand and Harvey Cohn. Remarks on principal factors in a relative cubic field. Journal of Number Theory, 3(2):226–239, May 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X71900400.

Stolarsky:1971:GBP

[80] Kenneth B. Stolarsky. Generalizations of the Bellavitis partition identities. Journal of Number Theory, 3(2):240–246, May 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X71900412.

Tietavainen:1971:PCS

[81] Aimo Tietäväinen. On a problem of Chowla and Shimura. Journal of Number Theory, 3(2):247–252, May 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X71900424.

Anonymous:1971:PA

[82] Anonymous. Pages 253–370 (August 1971). Journal of Number Theory, 3(3):??, August 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Schmidt:1971:LFA

[83] Wolfgang M. Schmidt. Linear forms with algebraic coefficients. I. Journal of Number Theory, 3(3):253–277, August 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X71900011.

Hall:1971:DDIb

[84] R. R. Hall. On the distribution of divisors of integers in residue classes (modk), III. Journal of Number Theory, 3(3):278–287, August 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X71900023.

Berndt:1971:AOCb

[85] Bruce C. Berndt. On the average order of a class of arithmetical functions, II. Journal of Number Theory, 3(3):288–305, August 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X71900035.

Brown:1971:SEO

[86] H. Brown and H. Zassenhaus. Some empirical observations on primitive roots. Journal of Number Theory, 3(3):306–309, August 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X71900047.

Fein:1971:RST

[87] Burton Fein, Basil Gordon, and John H. Smith. On the representation of −1 as a sum of two squares in an algebraic number field. Journal of Number Theory, 3(3):310–315, August 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X71900059.

Williams:1971:SS

[88] Kenneth S. Williams. On Salié’s sum. Journal of Number Theory, 3 (3):316–317, August 1971. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X71900060.

Furuta:1971:ZKR

[89] Yoshiomi Furuta. Über die zentrale Klassenzahl eines relativ-galoisschen Zahlkörpers. (German) [On the central class number of a relative Galois number field]. Journal of Number Theory, 3(3):318–322, August 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X71900072.

Goldstein:1971:GST

[90] Larry Joel Goldstein. A generalization of Stark’s theorem. Journal of Number Theory, 3(3):323–346, August 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X71900084.

Maurer:1971:ALR

[91] Donald Maurer. AF-lattices on a regular quadratic space. Journal of Number Theory, 3(3):347–363, August 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X71900096.

Brown:1971:GFS

[92] H. Brown and K. Mahler. A generalization of Farey sequences: Some exploration via the computer. Journal of Number Theory, 3(3):364–370, August 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X71900102.

Anonymous:1971:PN

[93] Anonymous. Pages 371–506 (November 1971). Journal of Number Theory, 3(4):??, November 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Pfeuffer:1971:EGT

[94] Horst Pfeuffer. Einklassige Geschlechter totalpositiver quadratischer Formen in totalreellen algebraischen Zahlkörpern. (German) [One-class relations of totally positive quadratic forms in totally real algebraic numbers]. Journal of Number Theory, 3(4):371–411, November 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X71900114.

Halter-Koch:1971:ATN

[95] Franz Halter-Koch. Arithmetische Theorie der Normalkörper von 2-Potenzgrad mit Diedergruppe. (German) [Arithmetic theory of normal bodies of 2-potency with dihedral group]. Journal of Number Theory, 3 (4):412–443, November 1971. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X71900126.

Draxl:1971:FAT

[96] P. K. J. Draxl. L-Funktionen Algebraischer Tori. (German) [L-functions algebraic tori]. Journal of Number Theory, 3(4):444–467, November 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X71900138.

Waid:1971:GAS

[97] Carter Waid. Generalized automorphs of the sum of four squares. Journal of Number Theory, 3(4):468–473, November 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X7190014X.

Orr:1971:RES

[98] Richard C. Orr. Remainder estimates for squarefree integers in arithmetic progression. Journal of Number Theory, 3(4):474–497, November 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X71900151.

Spira:1971:IRR

[99] Robert Spira. The integral representation for the Riemann Ξ-function. Journal of Number Theory, 3(4):498–501, November 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X71900163.

Levinson:1971:TBB

[100] N. Levinson. On theorems of Berlowitz and Berndt. Journal of Number Theory, 3(4):502–504, November 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X71900175.

Anonymous:1971:AI

[101] Anonymous. Author index. Journal of Number Theory, 3(4):505–506, November 1971. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X71900187.

Anonymous:1972:PF

[102] Anonymous. Pages 1–106 (February 1972). Journal of Number Theory, 4 (1):??, February 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1972:EB

[103] Anonymous. Editorial Board. Journal of Number Theory, 4(1):i, February 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X72900078.

Vaughan:1972:ASI

[104] R. C. Vaughan. On the addition of sequences of integers. Journal of Number Theory, 4(1):1–16, February 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X7290008X.

Wilson:1972:CDF

[105] Richard M. Wilson. Cyclotomy and difference families in elementary abelian groups. Journal of Number Theory, 4(1):17–47, February 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X72900091.

Bernstein:1972:DPJ

[106] Leon Bernstein. A 3-dimensional periodic Jacobi–Perron algorithm of period length 8. Journal of Number Theory, 4(1):48–69, February 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X72900108.

Halter-Koch:1972:EPR

[107] Franz Halter-Koch. Einseinheitengruppen und prime Restklassengruppen in quadratischen Zahlkörpern. (German) [Unit groups and prime remainder groups in quadratic number fields]. Journal of Number Theory, 4 (1):70–77, February 1972. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X7290011X.

Scharlau:1972:QRL

[108] Winfried Scharlau. Quadratic reciprocity laws. Journal of Number Theory, 4(1):78–97, February 1972. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X72900121.

Nijenhuis:1972:RIL

[109] Albert Nijenhuis and Herbert S. Wilf. Representations of integers by linear forms in nonnegative integers. Journal of Number Theory, 4(1):98–106, February 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X72900133.

Anonymous:1972:PAa

[110] Anonymous. Pages 107–223 (April 1972). Journal of Number Theory, 4 (2):??, April 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Ellison:1972:DE

[111] W. J. Ellison, F. Ellison, J. Pesek, C. E. Stahl, and D. S. Stall. The diophantine equation y^{2} + k = x^{3}. Journal of Number Theory, 4(2):107– 117, April 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X72900583.

Stepanov:1972:EPH

[112] S. A. Stepanov. An elementary proof of the Hasse–Weil theorem for hyperelliptic curves. Journal of Number Theory, 4(2):118–143, April 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X72900595.

Halter-Koch:1972:SGR

[113] Franz Halter-Koch. Ein Satz über die Geschlechter relativ-zyklischer Zahlkörper von Primzahlgrad und seine Anwendung auf biquadratischbizyklische Körper. (German) [A conjecture about the relations of relativecyclic number fields of prime number degree and its application to biquadratic–bicyclic fields]. Journal of Number Theory, 4(2):144–156, April 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X72900601.

Woods:1972:CSS

[114] A. C. Woods. Covering six space with spheres. Journal of Number Theory, 4(2):157–180, April 1972. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X72900613.

Kubota:1972:NCW

[115] K. K. Kubota. Note on a conjecture of W. Narkiewicz. Journal of Number Theory, 4(2):181–190, April 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X72900625.

Connors:1972:FCN

[116] Edward A. Connors. Finiteness of class number in characteristic 2. Journal of Number Theory, 4(2):191–218, April 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X72900637.

Gupta:1972:LJS

[117] Hansraj Gupta. The Legendre and Jacobi symbols for k-ic residues. Journal of Number Theory, 4(2):219–222, April 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X72900649.

Anonymous:1972:Ea

[118] Anonymous. Erratum. Journal of Number Theory, 4(2):223, April 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X72900650.

Anonymous:1972:PJ

[119] Anonymous. Pages 225–311 (June 1972). Journal of Number Theory, 4 (3):??, June 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Mahler:1972:JMJ

[120] K. Mahler. L. J. Mordell: 28 January, 1888 to 12 March, 1972. Journal of Number Theory, 4(3):iii–iv, June 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X72900480.

Grosswald:1972:RCV

[121] Emil Grosswald. Remarks concerning the values of the Riemann zeta function at integral, odd arguments. Journal of Number Theory, 4(3):225– 235, June 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X72900492.

Ennola:1972:RBC

[122] Veikko Ennola. On relations between cyclotomic units. Journal of Number Theory, 4(3):236–247, June 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X72900509.

Tillman:1972:BIQ

[123] Stephen J. Tillman. Binary integral quadratic forms over a certain class of function fields. Journal of Number Theory, 4(3):248–260, June 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X72900510.

Lindstrom:1972:SV

[124] Bernt Lindström. On B_{2}-sequences of vectors. Journal of Number Theory, 4(3):261–265, June 1972. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X72900522.

Camion:1972:ADS

[125] Paul Camion and H. B. Mann. Antisymmetric difference sets. Journal of Number Theory, 4(3):266–268, June 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X72900534.

Diderrich:1972:RP

[126] G. T. Diderrich and H. B. Mann. Representations by k-th powers in GF(q). Journal of Number Theory, 4(3):269–273, June 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X72900546.

Divis:1972:ALN

[127] Bohuslav Diviš. An analog to the Lagrange numbers. Journal of Number Theory, 4(3):274–285, June 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X72900558.

Goldstein:1972:IQF

[128] Larry Joel Goldstein. Imaginary quadratic fields of class number 2. Journal of Number Theory, 4(3):286–301, June 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X7290056X.

McEliece:1972:EPC

[129] Robert J. McEliece and Howard Rumsey. Euler products, cyclotomy, and coding. Journal of Number Theory, 4(3):302–311, June 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X72900571.

Anonymous:1972:PAb

[130] Anonymous. Pages 313–404 (August 1972). Journal of Number Theory, 4(4):??, August 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Hall:1972:PNT

[131] R. S. Hall. The prime number theorem for generalized primes. Journal of Number Theory, 4(4):313–320, August 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X72900662.

James:1972:RIQ

[132] D. G. James. Representations by integral quadratic forms. Journal of Number Theory, 4(4):321–329, August 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X72900674.

Dickson:1972:VRP

[133] T. J. Dickson. On Voronoi reduction of positive definite quadratic forms. Journal of Number Theory, 4(4):330–341, August 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X72900686.

Bleicher:1972:NAE

[134] M. N. Bleicher. A new algorithm for the expansion of Egyptian fractions. Journal of Number Theory, 4(4):342–382, August 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X72900698.

Ayoub:1972:GUC

[135] Christine W. Ayoub. On the group of units of certain rings. Journal of Number Theory, 4(4):383–403, August 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X72900704.

Anonymous:1972:Ec

[136] Anonymous. Erratum. Journal of Number Theory, 4(4):404, August 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X72900716.

Anonymous:1972:PO

[137] Anonymous. Pages 405–507 (October 1972). Journal of Number Theory, 4(5):??, October 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Mordell:1972:NPI

[138] L. J. Mordell. Note on the product of n inhomogeneous linear forms. Journal of Number Theory, 4(5):405–407, October 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X72900327.

Brown:1972:BQF

[139] Ezra Brown. Binary quadratic forms of determinant −pq. Journal of Number Theory, 4(5):408–410, October 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X72900339.

Sonn:1972:EPN

[140] Jack Sonn. On the embedding problem for nonsolvable Galois groups of algebraic number fields: Reduction theorems. Journal of Number Theory, 4(5):411–436, October 1972. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X72900340.

Huxley:1972:ISS

[141] M. N. Huxley. Irregularity in sifted sequences. Journal of Number Theory, 4(5):437–454, October 1972. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X72900352.

Finkelstein:1972:TNW

[142] Raphael Finkelstein and Hymie London. On triangular numbers which are sums of consecutive squares. Journal of Number Theory, 4(5):455–462, October 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X72900364.

Steinig:1972:ICA

[143] J. Steinig. On an integral connected with the average order of a class of arithmetical functions. Journal of Number Theory, 4(5):463–468, October 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X72900376.

Nymann:1972:PPI

[144] J. E. Nymann. On the probability that k positive integers are relatively prime. Journal of Number Theory, 4(5):469–473, October 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X72900388.

Barnes:1972:SAN

[145] F. W. Barnes. On the stufe of an algebraic number field. Journal of Number Theory, 4(5):474–476, October 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X7290039X.

Niederreiter:1972:CSL

[146] Harald Niederreiter. On a class of sequences of lattice points. Journal of Number Theory, 4(5):477–502, October 1972. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X72900406.

Morikawa:1972:URQ

[147] Ryozo Morikawa. On units of real quadratic fields. Journal of Number Theory, 4(5):503–507, October 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X72900418.

Anonymous:1972:PD

[148] Anonymous. Pages 509–597 (December 1972). Journal of Number Theory, 4(6):??, December 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Elliott:1972:DFQ

[149] P. D. T. A. Elliott. On distribution functions (mod1): quantitative Fourier inversion. Journal of Number Theory, 4(6):509–522, December 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X72900248.

Wolke:1972:MCN

[150] Dieter Wolke. Moments of class numbers, III. Journal of Number Theory, 4 (6):523–531, December 1972. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X7290025X.

Dressler:1972:EPT

[151] Robert E. Dressler. An elementary proof of a theorem of Erdős on the sum of divisors function. Journal of Number Theory, 4(6):532–536, December 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X72900261.

Shanks:1972:NTQ

[152] Daniel Shanks. New types of quadratic fields having three invariants divisible by 3. Journal of Number Theory, 4(6):537–556, December 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X72900273.

Garbanati:1972:SCQ

[153] Dennis Garbanati and Robert C. Thompson. Skew circulant quadratic forms. Journal of Number Theory, 4(6):557–572, December 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X72900285.

Iyanaga:1972:AHF

[154] Kenichi Iyanaga. Arithmetic of Hermitian forms and related forms. Journal of Number Theory, 4(6):573–586, December 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X72900297.

Kubota:1972:FPU

[155] K. K. Kubota. Factors of polynomials under composition. Journal of Number Theory, 4(6):587–595, December 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X72900303.

Anonymous:1972:AI

[156] Anonymous. Author index. Journal of Number Theory, 4(6):596–597, December 1972. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X72900315.

Anonymous:1973:PF

[157] Anonymous. Pages 1–97 (February 1973). Journal of Number Theory, 5 (1):??, February 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1973:EB

[158] Anonymous. Editorial Board. Journal of Number Theory, 5(1):i, February 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X73900528.

France:1973:SSV

[159] Michel Mendes France. Les suites à spectre vide et la répartition modulo 1. (French) [Empty spectrum suites and distribution modulo 1]. Journal of Number Theory, 5(1):1–15, February 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X7390053X.

Matthews:1973:HFL

[160] K. R. Matthews. Hermitian forms and the large and small sieves. Journal of Number Theory, 5(1):16–23, February 1973. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X73900541.

Bertness:1973:CRS

[161] Charles H. Bertness and Leon R. McCulloh. Corresponding residue systems in nonnormal extensions of prime degree. Journal of Number Theory, 5(1):24–42, February 1973. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X73900553.

Lehmer:1973:SPC

[162] D. H. Lehmer. Some properties of circulants. Journal of Number Theory, 5(1):43–54, February 1973. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X73900565.

Onishi:1973:TTD

[163] H. Onishi. A Tauberian theorem on Dirichlet series. Journal of Number Theory, 5(1):55–57, February 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X73900577.

Goldstein:1973:CNC

[164] Larry Joel Goldstein. On the class numbers of cyclotomic fields. Journal of Number Theory, 5(1):58–63, February 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X73900589.

Vaughan:1973:SAM

[165] R. C. Vaughan. Some applications of Montgomery’s sieve. Journal of Number Theory, 5(1):64–79, February 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X73900590.

Tijdeman:1973:ART

[166] R. Tijdeman. An auxiliary result in the theory of transcendental numbers. Journal of Number Theory, 5(1):80–94, February 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X73900607.

Waterhouse:1973:PEC

[167] William C. Waterhouse. Pieces of eight in class groups of quadratic fields. Journal of Number Theory, 5(1):95–97, February 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X73900619.

Anonymous:1973:PAa

[168] Anonymous. Pages 99–178 (April 1973). Journal of Number Theory, 5 (2):??, April 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Benson:1973:CUC

[169] C. T. Benson and B. T. Weber. Computing units in certain orders of algebraic integers. Journal of Number Theory, 5(2):99–107, April 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X73900620.

Evans:1973:FRH

[170] Ronald Evans. A fundamental region for Hecke’s modular group. Journal of Number Theory, 5(2):108–115, April 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X73900632.

Stout:1973:CRS

[171] William T. Stout. Corresponding residue systems in normal extensions. Journal of Number Theory, 5(2):116–122, April 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X73900644.

Schweiger:1973:SRK

[172] F. Schweiger and M. Waterman. Some remarks on Kuzmin’s theorem for F -expansions. Journal of Number Theory, 5(2):123–131, April 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X73900656.

Camion:1973:PR

[173] Paul Camion, L. S. Levy, and H. B. Mann. Prüfer rings. Journal of Number Theory, 5(2):132–138, April 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X73900668.

Moser:1973:RCS

[174] Claude Moser. Représentation de −1 comme somme de carrés dans un corps cyclotomique quelconque. (French) [Representation of −1 as sum of squares in any cyclotomic body]. Journal of Number Theory, 5(2):139– 141, April 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X7390067X.

Niemeier:1973:DQF

[175] Hans-Volker Niemeier. Definite quadratische Formen der Dimension 24 und Diskriminante 1. (German) [Definite quadratic forms of dimension 24 and discriminant 1]. Journal of Number Theory, 5(2):142–178, April 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X73900681.

Anonymous:1973:PJ

[176] Anonymous. Pages 179–251 (June 1973). Journal of Number Theory, 5 (3):??, June 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Yamada:1973:SCG

[177] Toshihiko Yamada. Simple components of group algebras Q[G] central over real quadratic fields. Journal of Number Theory, 5(3):179–190, June 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X73900437.

Waldschmidt:1973:SDH

[178] Michel Waldschmidt. Solution du huitième problème de Schneider. (French) [Solution of Schneider’s eighth problem]. Journal of Number Theory, 5(3):191–202, June 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X73900449.

Jutila:1973:CSC

[179] Matti Jutila. On character sums and class numbers. Journal of Number Theory, 5(3):203–214, June 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X73900450.

Hooley:1973:IBN

[180] C. Hooley. On the intervals between numbers that are sums of two squares: II. Journal of Number Theory, 5(3):215–217, June 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X73900462.

Golomb:1973:NAF

[181] Solomon W. Golomb. A new arithmetic function of combinatorial significance. Journal of Number Theory, 5(3):218–223, June 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X73900474.

Baker:1973:PC

[182] A. Baker, B. J. Birch, and E. A. Wirsing. On a problem of Chowla. Journal of Number Theory, 5(3):224–236, June 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X73900486.

Weinberger:1973:RQF

[183] P. J. Weinberger. Real quadratic fields with class numbers divisible by n. Journal of Number Theory, 5(3):237–241, June 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X73900498.

vanderWaall:1973:P

[184] R. W. van der Waall. On a property of tan x. Journal of Number Theory, 5(3):242–244, June 1973. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X73900504.

Sweet:1973:TDA

[185] Melvin M. Sweet. A theorem in diophantine approximations. Journal of Number Theory, 5(3):245–251, June 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X73900516.

Anonymous:1973:PAb

[186] Anonymous. Pages 253–327 (August 1973). Journal of Number Theory, 5(4):??, August 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Adachi:1973:GKC

[187] Norio Adachi. Generalization of Kummer’s criterion for divisibility of class numbers. Journal of Number Theory, 5(4):253–265, August 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X73900693.

Parry:1973:PRB

[188] Charles J. Parry. Primes represented by binary quadratic forms. Journal of Number Theory, 5(4):266–270, August 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X7390070X.

Brandler:1973:RPR

[189] Jacob A. Brandler. Residuacity properties of real quadratic units. Journal of Number Theory, 5(4):271–286, August 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X73900711.

Dodson:1973:FDC

[190] M. M. Dodson. On a function due to S. Chowla. Journal of Number Theory, 5(4):287–292, August 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X73900723.

Yap:1973:MSF

[191] Hian-Poh Yap. Maximal sum-free sets in finite Abelian groups, III. Journal of Number Theory, 5(4):293–300, August 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X73900735.

Robinson:1973:NPA

[192] Abraham Robinson. Nonstandard points on algebraic curves. Journal of Number Theory, 5(4):301–327, August 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X73900747.

Anonymous:1973:PO

[193] Anonymous. Pages 329–419 (October 1973). Journal of Number Theory, 5(5):??, October 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Pall:1973:RQF

[194] Gordon Pall. Report of the quadratic forms conference: Louisiana State University, Baton Rouge, Louisiana 70803. Journal of Number Theory, 5 (5):329–330, October 1973. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X73900322.

Pall:1973:MMP

[195] Gordon Pall. A method of making proofs involving numerous cases readable. Journal of Number Theory, 5(5):331, October 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X73900334.

Gerstein:1973:OSC

[196] Larry J. Gerstein. Orthogonal splitting and class numbers of quadratic forms. Journal of Number Theory, 5(5):332–338, October 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X73900346.

Hsia:1973:WRS

[197] J. S. Hsia. On the Witt ring and some arithmetical invariants of quadratic forms. Journal of Number Theory, 5(5):339–355, October 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X73900358.

Hsia:1973:RGQ

[198] J. S. Hsia and Robert P. Johnson. Round and group quadratic forms over global fields. Journal of Number Theory, 5(5):356–366, October 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X7390036X.

Elman:1973:PFT

[199] Richard Elman and T. Y. Lam. Pfister forms and their applications. Journal of Number Theory, 5(5):367–378, October 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X73900371.

Maurer:1973:TFA

[200] Donald Maurer. The trace-form of an algebraic number field. Journal of Number Theory, 5(5):379–384, October 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X73900383.

Lehmer:1973:CCQ

[201] Emma Lehmer. On the cubic character of quadratic units. Journal of Number Theory, 5(5):385–389, October 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X73900395.

Conway:1973:IQF

[202] J. H. Conway. Invariants for quadratic forms. Journal of Number Theory, 5(5):390–404, October 1973. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X73900401.

Carlitz:1973:BMS

[203] Leonard Carlitz. Bulygin’s method for sums of squares. Journal of Number Theory, 5(5):405–412, October 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X73900413.

Brown:1973:PDC

[204] Ezra Brown. The power of 2 dividing the class-number of a binary quadratic discriminant. Journal of Number Theory, 5(5):413–419, October 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X73900425.

Anonymous:1973:PD

[205] Anonymous. Pages 421–550 (December 1973). Journal of Number Theory, 5(6):??, December 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Estes:1973:SGB

[206] Dennis R. Estes and Gordon Pall. Spinor genera of binary quadratic forms. Journal of Number Theory, 5(6):421–432, December 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X73900127.

Estes:1973:RLJ

[207] Dennis R. Estes and Gordon Pall. A reconsideration of Legendre– Jacobi symbols. Journal of Number Theory, 5(6):433–434, December 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X73900139.

Pall:1973:BQC

[208] Gordon Pall. Binary quadratic and cubic forms, and unipositive matrices. Journal of Number Theory, 5(6):435–443, December 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X73900140.

James:1973:SOG

[209] D. G. James. The structure of orthogonal groups of 2-adic unimodular quadratic forms. Journal of Number Theory, 5(6):444–455, December 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X73900152.

Chang:1973:UGI

[210] Chan nan Chang. The unitary group over the integers of a quaternion algebra. Journal of Number Theory, 5(6):456–476, December 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X73900164.

Connors:1973:AOG

[211] Edward A. Connors. Automorphism of orthogonal groups in characteristic 2. Journal of Number Theory, 5(6):477–501, December 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X73900176.

Feaux:1973:DMC

[212] Charles Feaux. Divisor matrices in the Cayley ring. Journal of Number Theory, 5(6):502–523, December 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X73900188.

Rice:1973:RIQ

[213] Bart F. Rice. Rings of integral quaternions. Journal of Number Theory, 5 (6):524–536, December 1973. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X7390019X.

Cordes:1973:SOV

[214] Craig M. Cordes. Sequences of orthogonal vectors with a fixed norm. Journal of Number Theory, 5(6):537–540, December 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X73900206.

Caron:1973:UQF

[215] Richard Caron. Unimodular quadratic forms via skew orthogonal matrices. Journal of Number Theory, 5(6):541, December 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X73900218.

Cheema:1973:PQF

[216] M. S. Cheema and G. Pall. Positive quadratic forms of small determinants. Journal of Number Theory, 5(6):542, December 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X7390022X.

Couvillon:1973:PNM

[217] Lawrence A. Couvillon. Partitionings and nonpartitionings by maximal isotropic spaces. Journal of Number Theory, 5(6):543, December 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X73900231.

Johnson:1973:AUG

[218] Arnold A. Johnson. The automorphisms of the unitary groups over infinite fields. Journal of Number Theory, 5(6):543, December 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X73900243.

Jones:1973:QGQ

[219] Burton W. Jones. Quasi-genera of quadratic forms. Journal of Number Theory, 5(6):544, December 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X73900255.

Lehmer:1973:EQF

[220] D. H. Lehmer. The economics of quadratic form calculations. Journal of Number Theory, 5(6):544–545, December 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X73900267.

Matherne:1973:SCG

[221] John W. Matherne. The structure of the class group in orders of binary quadratic forms. Journal of Number Theory, 5(6):545–546, December 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X73900279.

Pall:1973:GCN

[222] Gordon Pall and Charles Parry. Generalizations of a class-number formula of Gauss. Journal of Number Theory, 5(6):546–547, December 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X73900280.

Riehm:1973:EBF

[223] Carl Riehm. The equivalence of bilinear forms. Journal of Number Theory, 5(6):547, December 1973. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X73900292.

Taussky:1973:FIM

[224] Olga Taussky. The factorization of an integral matrix into a product of two integral symmetric matrices. I. Journal of Number Theory, 5(6):547–548, December 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X73900309.

Anonymous:1973:AI

[225] Anonymous. Author index. Journal of Number Theory, 5(6):549–550, December 1973. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X73900310.

Anonymous:1974:PF

[226] Anonymous. Pages 1–71 (February 1974). Journal of Number Theory, 6 (1):??, February 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1974:EB

[227] Anonymous. Editorial Board. Journal of Number Theory, 6(1):i, February 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X74900018.

An:1974:ACC

[228] Chungming An. On the analytic continuation of a certain Dirichlet series. Journal of Number Theory, 6(1):1–6, February 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X7490002X.

Newman:1974:GTC

[229] D. J. Newman. A gradually turning curve must come near a lattice point. Journal of Number Theory, 6(1):7–10, February 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X74900031.

Brownawell:1974:SDA

[230] W. Dale Brownawell. Sequences of Diophantine approximations. Journal of Number Theory, 6(1):11–21, February 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X74900043.

Brownawell:1974:AIC

[231] W. Dale Brownawell. The algebraic independence of certain numbers related by the exponential function. Journal of Number Theory, 6(1):22– 31, February 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X74900055.

Ullom:1974:IRA

[232] S. Ullom. Integral representations afforded by ambiguous ideals in some abelian extensions. Journal of Number Theory, 6(1):32–49, February 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X74900067.

Heisler:1974:DFF

[233] Brother Joseph Heisler. Diagonal forms over finite fields. Journal of Number Theory, 6(1):50–51, February 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X74900079.

Erdos:1974:DVC

[234] P. Erdős and R. R. Hall. On the distribution of values of certain divisor functions. Journal of Number Theory, 6(1):52–63, February 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X74900080.

Taussky:1974:RCC

[235] Olga Taussky. A result concerning classes of matrices. Journal of Number Theory, 6(1):64–71, February 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X74900092.

Anonymous:1974:PAa

[236] Anonymous. Pages 73–159 (April 1974). Journal of Number Theory, 6 (2):??, April 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Bantegnie:1974:CCF

[237] Robert Bantegnie. La “conjecture des cylindres”. (French) [The conjecture of cylinders]. Journal of Number Theory, 6(2):73–98, April 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X74900468.

Ishida:1974:CNA

[238] Makoto Ishida. Class numbers of algebraic number fields of Eisenstein type, II. Journal of Number Theory, 6(2):99–104, April 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X7490047X.

Choi:1974:EPN

[239] S. L. G. Choi. On an extremal problem in number theory. Journal of Number Theory, 6(2):105–111, April 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X74900481.

Subbarao:1974:OEF

[240] M. V. Subbarao and D. Suryanarayana. On the order of the error function of the (k, r)-integers. Journal of Number Theory, 6(2):112–123, April 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X74900493.

Sehgal:1974:PSG

[241] Surinder K. Sehgal. F_{Π} projectors of Π soluble groups. Journal of Number Theory, 6(2):124–127, April 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X7490050X.

Swinnerton-Dyer:1974:NLP

[242] H. P. F. Swinnerton-Dyer. The number of lattice points on a convex curve. Journal of Number Theory, 6(2):128–135, April 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X74900511.

Chowla:1974:CPC

[243] S. Chowla and P. Hartung. Congruence properties of class numbers of quadratic fields. Journal of Number Theory, 6(2):136–137, April 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X74900523.

Sonn:1974:LEP

[244] Jack Sonn. Localizability of the embedding problem with symplectic kernel. Journal of Number Theory, 6(2):138–157, April 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X74900535.

Chowla:1974:NCS

[245] Paromita Chowla. On the nonvanishing of a certain L-series at s = 12. Journal of Number Theory, 6(2):158–159, April 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X74900547.

Anonymous:1974:PJ

[246] Anonymous. Pages 161–237 (June 1974). Journal of Number Theory, 6 (3):??, June 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Delange:1974:IVA

[247] Hubert Delange. On integral-valued additive functions, II. Journal of Number Theory, 6(3):161–170, June 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X74900109.

Pandya:1974:AFS

[248] Gautam N. Pandya. On automorphisms of finite simple Chevalley groups. Journal of Number Theory, 6(3):171–184, June 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X74900110.

Brown:1974:CNC

[249] Ezra Brown. Class numbers of complex quadratic fields. Journal of Number Theory, 6(3):185–191, June 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X74900122.

Hawkins:1974:RSI

[250] David Hawkins. Random sieves, II. Journal of Number Theory, 6(3):192– 200, June 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X74900134.

Guthery:1974:TTO

[251] Scott B. Guthery. A transformation theorem for one-dimensional F - expansions. Journal of Number Theory, 6(3):201–210, June 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X74900146.

Fried:1974:HIT

[252] Michael Fried. On Hilbert’s Irreducibility Theorem. Journal of Number Theory, 6(3):211–231, June 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X74900158.

Coleman:1974:EWR

[253] D. B. Coleman. Embedding Witt rings of Dedekind domains. Journal of Number Theory, 6(3):232–237, June 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X7490016X.

Anonymous:1974:PAb

[254] Anonymous. Pages 239–333 (August 1974). Journal of Number Theory, 6(4):??, August 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Pandya:1974:AGA

[255] Gautam N. Pandya. Algebraic groups and automorphisms of finite simple Chevalley groups. Journal of Number Theory, 6(4):239–247, August 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X74900171.

Shorey:1974:SAN

[256] T. N. Shorey. On the sum ∑_{k=1}^{3} |2^{π^{k}} − α_{k}|, α_{k} algebraic numbers. Journal of Number Theory, 6(4):248–260, August 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X74900183.

Chowla:1974:NHH

[257] S. Chowla and M. J. DeLeon. A note on the Hecke hypothesis and the determination of imaginary quadratic fields with class-number 1. Journal of Number Theory, 6(4):261–263, August 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X74900195.

Bernstein:1974:ZF

[258] Leon Bernstein. Zeros of the functions f(n) = ∑_{i=0}(−1)^{i}(i^{n−2i}). Journal of Number Theory, 6(4):264–270, August 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X74900201.

Chowla:1974:NHA

[259] S. Chowla and P. Hartung. A note on the hypothesis that L(s, χ) > 0 for all real non-principal characters χ and for all s > 0. Journal of Number Theory, 6(4):271–275, August 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X74900213.

Hartung:1974:PEI

[260] P. Hartung. Proof of the existence of infinitely many imaginary quadratic fields whose class number is not divisible by 3. Journal of Number Theory, 6(4):276–278, August 1974. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X74900225.

Hartung:1974:ECC

[261] P. Hartung. Explicit construction of a class of infinitely many imaginary quadratic fields whose class number is divisible by 3. Journal of Number Theory, 6(4):279–281, August 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X74900237.

Risman:1974:NPT

[262] Lawrence J. Risman. A new proof of the three squares theorem. Journal of Number Theory, 6(4):282–283, August 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X74900249.

Schanuel:1974:ECT

[263] Stephen H. Schanuel. An extension of Chevalley’s theorem to congruences modulo prime powers. Journal of Number Theory, 6(4):284–290, August 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X74900250.

Ponomarev:1974:CND

[264] Paul Ponomarev. Class numbers of definite quaternary forms with square discriminant. Journal of Number Theory, 6(4):291–317, August 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X74900262.

Lakein:1974:BCP

[265] Richard B. Lakein. Bounds for consecutive k-th power residues in the Eisenstein integers. Journal of Number Theory, 6(4):318–323, August 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X74900274.

Nathanson:1974:MBM

[266] Melvyn B. Nathanson. Minimal bases and maximal nonbases in additive number theory. Journal of Number Theory, 6(4):324–333, August 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X74900286.

Anonymous:1974:PO

[267] Anonymous. Pages 335–403 (October 1974). Journal of Number Theory, 6(5):??, October 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Leonard:1974:FGP

[268] Philip A. Leonard. Factorizations of general polynomials. Journal of Number Theory, 6(5):335–338, October 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X74900298.

Rogers:1974:LTQ

[269] Kenneth Rogers. Legendre’s Theorem and quadratic reciprocity. Journal of Number Theory, 6(5):339–344, October 1974. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X74900304.

Singmaster:1974:PF

[270] David Singmaster. On polynomial functions ( mod m). Journal of Number Theory, 6(5):345–352, October 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X74900316.

Williams:1974:MTA

[271] Kenneth S. Williams. Merten’s theorem for arithmetic progressions. Journal of Number Theory, 6(5):353–359, October 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X74900328.

Chang:1974:NSC

[272] P. L. Chang. On the number of square classes of a field with finite stufe. Journal of Number Theory, 6(5):360–368, October 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X7490033X.

Gold:1974:NCE

[273] Robert Gold. The nontriviality of certain Z_{l}-extensions. Journal of Number Theory, 6(5):369–373, October 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X74900341.

Osgood:1974:SDA

[274] Charles F. Osgood. The simultaneous diophantine approximation of values of functions satisfying certain n variable q-difference equations. Journal of Number Theory, 6(5):374–400, October 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X74900353.

Brudno:1974:ESS

[275] Simcha Brudno and Irving Kaplansky. Equal sums of sixth powers. Journal of Number Theory, 6(5):401–403, October 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X74900365.

Anonymous:1974:PD

[276] Anonymous. Pages 405–487 (December 1974). Journal of Number Theory, 6(6):??, December 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Turan:1974:PP

[277] P. Turán. On a property of partitions. Journal of Number Theory, 6 (6):405–411, December 1974. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X74900377.

Bojanic:1974:SPM

[278] R. Bojanic. A simple proof of Mahler’s theorem on approximation of continuous functions of a p-adic variable by polynomials. Journal of Number Theory, 6(6):412–415, December 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X74900389.

Choi:1974:PI

[279] S. L. G. Choi and P. Erdős. On products of integers. Journal of Number Theory, 6(6):416–421, December 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X74900390.

Bambah:1974:TD

[280] R. P. Bambah and A. C. Woods. On a theorem of Dyson. Journal of Number Theory, 6(6):422–433, December 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X74900407.

Szekeres:1974:ND

[281] George Szekeres. On the number of divisors of x^{2}+x+A. Journal of Number Theory, 6(6):434–442, December 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X74900419.

Ankeny:1974:URF

[282] N. C. Ankeny and J. S. Hsia. Unimodular round forms over dyadic local fields. Journal of Number Theory, 6(6):443–447, December 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X74900420.

Schmidt:1974:LBN

[283] Wolfgang M. Schmidt. A lower bound for the number of solutions of equations over finite fields. Journal of Number Theory, 6(6):448–480, December 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X74900432.

Sprindzuk:1974:RNN

[284] V. G. Sprindžuk. Representation of numbers by the norm forms with two dominating variables. Journal of Number Theory, 6(6):481–486, December 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X74900444.

Anonymous:1974:AI

[285] Anonymous. Author index. Journal of Number Theory, 6(6):487, December 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X74900456.

Anonymous:1975:PF

[286] Anonymous. Pages 1–120 (February 1975). Journal of Number Theory, 7 (1):??, February 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1975:EB

[287] Anonymous. Editorial Board. Journal of Number Theory, 7(1):i, February 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X75900013.

Leopoldt:1975:SKN

[288] H. W. Leopoldt. Sigekatu Kuroda: November 11, 1905–November 3, 1972. Journal of Number Theory, 7(1):1–4, February 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X75900025.

Alter:1975:DER

[289] Ronald Alter and K. K. Kubota. The Diophantine equation x^{2} + 11 = 3^{n} and a related sequence. Journal of Number Theory, 7(1):5–10, February 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X75900037.

Leitzel:1975:AFF

[290] James R. C. Leitzel, Manohar L. Madan, and Clifford S. Queen. Algebraic function fields with small class number. Journal of Number Theory, 7 (1):11–27, February 1975. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X75900049.

Jogdeo:1975:MAR

[291] Kumar Jogdeo and W. Molenaar. Δ-Monotone arrangements of real numbers. Journal of Number Theory, 7(1):28–32, February 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X75900050.

Diderrich:1975:ATA

[292] George T. Diderrich. An addition theorem for Abelian groups of order pq. Journal of Number Theory, 7(1):33–48, February 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X75900062.

Rehm:1975:GSR

[293] Hans Peter Rehm. Über die gruppentheoretische Struktur der Relationen zwischen Relativnormabbildungen in endlichen Galoisschen Körpererweiterungen. (German) [On the group-theoretical structure of the relations between relative norm maps in finite Galoisian field extensions]. Journal of Number Theory, 7(1):49–70, February 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X75900074.

Yokoi:1975:DIP

[294] Hideo Yokoi. On the distribution of irregular primes. Journal of Number Theory, 7(1):71–76, February 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X75900086.

Garbanati:1975:AGM

[295] Dennis Garbanati. Abelian group matrices over the p-adic and rational integers. Journal of Number Theory, 7(1):77–90, February 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X75900098.

Carlson:1975:GSI

[296] David Carlson. Good sequences of integers. Journal of Number Theory, 7(1):91–104, February 1975. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X75900104.

Ayoub:1975:TC

[297] Raymond Ayoub. On a theorem of S. Chowla. Journal of Number Theory, 7(1):105–107, February 1975. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X75900116.

Ayoub:1975:TI

[298] Raymond Ayoub. On a theorem of Iwasawa. Journal of Number Theory, 7(1):108–120, February 1975. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X75900128.

Anonymous:1975:PM

[299] Anonymous. Pages 121–265 (May 1975). Journal of Number Theory, 7 (2):??, May 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Robinson:1975:FTS

[300] A. Robinson and P. Roquette. On the finiteness theorem of Siegel and Mahler concerning diophantine equations. Journal of Number Theory, 7(2):121–176, May 1975. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X7590013X.

Goldfeld:1975:DSI

[301] D. Goldfeld. On Dirichlet series with identical beginnings. Journal of Number Theory, 7(2):177–183, May 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X75900141.

Szalay:1975:DPR

[302] Michael Szalay. On the distribution of the primitive roots of a prime. Journal of Number Theory, 7(2):184–188, May 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X75900153.

Hazlewood:1975:SPI

[303] D. G. Hazlewood. Sums over positive integers with few prime factors. Journal of Number Theory, 7(2):189–207, May 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X75900165.

Feaux:1975:FCC

[304] C. J. Feaux and J. T. Hardy. Factorization in certain Cayley rings. Journal of Number Theory, 7(2):208–220, May 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X75900177.

Herzberg:1975:IS

[305] Norman P. Herzberg. Integer solutions of by^{2} + p^{n} = x^{3}. Journal of Number Theory, 7(2):221–234, May 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X75900189.

Stender:1975:FGR

[306] Hans-Joachim Stender. Eine Formel für Grundeinheiten in reinen algebraischen Zahlkörpern dritten, vierten und sechsten Grades. (German) [A formula for basic units in pure algebraic number fields of third, fourth and sixth degree]. Journal of Number Theory, 7(2):235–250, May 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X75900190.

Johnson:1975:APC

[307] Wells Johnson. p-adic Proofs of congruences for the Bernoulli numbers. Journal of Number Theory, 7(2):251–265, May 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X75900207.

Anonymous:1975:PA

[308] Anonymous. Pages 267–356 (August 1975). Journal of Number Theory, 7(3):??, August 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Jones:1975:COIa

[309] Antony J. Jones. Cyclic overlattices (III). Journal of Number Theory, 7 (3):267–282, August 1975. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X75900219.

Jones:1975:COIb

[310] Antony J. Jones. Cyclic overlattices (IV). Journal of Number Theory, 7 (3):283–292, August 1975. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X75900220.

Hightower:1975:ACN

[311] Collin J. Hightower. Approximation to Complex numbers by certain biquadratic numbers. Journal of Number Theory, 7(3):293–309, August 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X75900232.

Pleasants:1975:CPA

[312] P. A. B. Pleasants. Cubic polynomials over algebraic number fields. Journal of Number Theory, 7(3):310–344, August 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X75900244.

Janusz:1975:SGC

[313] G. J. Janusz. The Schur group of cyclotomic fields. Journal of Number Theory, 7(3):345–352, August 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X75900256.

Tietavainen:1975:PCC

[314] Aimo Tietäväinen. Proof of a conjecture of S. Chowla. Journal of Number Theory, 7(3):353–356, August 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X75900268.

Anonymous:1975:PN

[315] Anonymous. Pages 357–446 (November 1975). Journal of Number Theory, 7(4):??, November 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Zajta:1975:SDE

[316] Aurel J. Zajta. On the solution of the diophantine equation A^{3} + B^{3} + C^{3} + D^{3} = 0. Journal of Number Theory, 7(4):357–360, November 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X75900372.

Williams:1975:NC

[317] Kenneth S. Williams. Note on cubics over GF(2^{n}) and GF(3^{n}). Journal of Number Theory, 7(4):361–365, November 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X75900384.

Wou:1975:SSA

[318] Y. F. Wou. Sums of sets in the Abelian group of type (5, 5). Journal of Number Theory, 7(4):366–370, November 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X75900396.

MacCluer:1975:UM

[319] C. R. MacCluer and Charles J. Parry. Units of modulus 1. Journal of Number Theory, 7(4):371–375, November 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X75900402.

Parmenter:1975:NAG

[320] M. M. Parmenter and S. K. Sehgal. Non-Archimedean group algebras. Journal of Number Theory, 7(4):376–384, November 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X75900414.

Parry:1975:UAN

[321] Charles J. Parry. Units of algebraic numberfields. Journal of Number Theory, 7(4):385–388, November 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X75900426. See correction [413].

Richmond:1975:ARP

[322] L. B. Richmond. Asymptotic relations for partitions. Journal of Number Theory, 7(4):389–405, November 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X75900438.

Nymann:1975:PPI

[323] J. E. Nymann. On the probability that k positive integers are relatively prime II. Journal of Number Theory, 7(4):406–412, November 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X7590044X.

Berndt:1975:CAP

[324] Bruce C. Berndt. Character analogues of the Poisson and Euler– MacLaurin summation formulas with applications. Journal of Number Theory, 7(4):413–445, November 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X75900451.

Anonymous:1975:AI

[325] Anonymous. Author index. Journal of Number Theory, 7(4):446, November 1975. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X75900463.

Anonymous:1976:PF

[326] Anonymous. Pages 1–130 (February 1976). Journal of Number Theory, 8 (1):??, February 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1976:EB

[327] Anonymous. Editorial Board. Journal of Number Theory, 8(1):i, February 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X76900160.

Whiteman:1976:HMO

[328] Albert Leon Whiteman. Hadamard matrices of order 4(2p+1). Journal of Number Theory, 8(1):1–11, February 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X76900172.

Novak:1976:LPM

[329] Břetislav Novák. On lattice points in multidimensional ellipsoids: Problem of centers. Journal of Number Theory, 8(1):12–39, February 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X76900184.

Straus:1976:DR

[330] E. G. Straus. Differences of residues (modp). Journal of Number Theory, 8(1):40–42, February 1976. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X76900196.

Jordan:1976:LDG

[331] J. H. Jordan and J. R. Rabung. Local distribution of Gaussian primes. Journal of Number Theory, 8(1):43–51, February 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X76900202.

Olson:1976:CPE

[332] John E. Olson. On a combinatorial problem of Erdős, Ginzburg, and Ziv. Journal of Number Theory, 8(1):52–57, February 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X76900214.

Gaddis:1976:CFI

[333] T. A. Gaddis. G-Closed fields and imbeddings of quadratic number fields. Journal of Number Theory, 8(1):58–72, February 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X76900226.

Sehgal:1976:ECS

[334] Surinder K. Sehgal. On the existence of Cartan subgroups of finite groups. Journal of Number Theory, 8(1):73–83, February 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X76900238.

Gerth:1976:CGC

[335] Frank Gerth. On 3-class groups of cyclic cubic extensions of certain number fields. Journal of Number Theory, 8(1):84–98, February 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X7690024X.

Weidner:1976:PLL

[336] Hans G. Weidner. A periodicity lemma in linear diophantine analysis. Journal of Number Theory, 8(1):99–108, February 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X76900251.

Johnsen:1976:CSL

[337] E. C. Johnsen and T. Storer. Combinatorial structures in loops, III. Difference sets in special cyclic neofields. Journal of Number Theory, 8(1): 109–130, February 1976. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X76900263.

Anonymous:1976:PM

[338] Anonymous. Pages 131–250 (May 1976). Journal of Number Theory, 8 (2):??, May 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Gerig:1976:SPP

[339] S. Gerig. A simple proof of the Prime Number Theorem. Journal of Number Theory, 8(2):131–136, May 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X76900962.

Storer:1976:MG

[340] Thomas Storer. Modular graphs. Journal of Number Theory, 8(2):137– 156, May 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X76900974.

Bleicher:1976:DEF

[341] M. N. Bleicher and P. Erdős. Denominators of Egyptian fractions. Journal of Number Theory, 8(2):157–168, May 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X76900986.

Pandya:1976:SGE

[342] Gautam N. Pandya. Splitting of group extensions. Journal of Number Theory, 8(2):169–183, May 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X76900998.

Pizer:1976:PCN

[343] Arnold Pizer. On the 2-part of the class number of imaginary quadratic number fields. Journal of Number Theory, 8(2):184–192, May 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X76901001.

Weinstein:1976:SCP

[344] G. Weinstein. Some covering and packing results in number theory. Journal of Number Theory, 8(2):193–205, May 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X76901013.

Cohn:1976:FFR

[345] Harvey Cohn. Finiteness of the formal ring of a quadratic sector. Journal of Number Theory, 8(2):206–217, May 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X76901025.

Benkoski:1976:PPI

[346] Stanley J. Benkoski. The probability that k positive integers are relatively r-prime. Journal of Number Theory, 8(2):218–223, May 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X76901037.

Christie:1976:PDR

[347] M. R. Christie. Positive definite rational functions of two variables which are not the sum of three squares. Journal of Number Theory, 8(2):224– 232, May 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X76901049.

Karian:1976:CPF

[348] Z. A. Karian. The covering problem for the fields Q(−1)^{12} and Q(−3)^{12}. Journal of Number Theory, 8(2):233–244, May 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X76901050.

Washington:1976:NAF

[349] L. C. Washington. A note on p-adic L-functions. Journal of Number Theory, 8(2):245–250, May 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X76901062.

Anonymous:1976:PA

[350] Anonymous. Pages 251–368 (August 1976). Journal of Number Theory, 8(3):??, August 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Burshtein:1976:NEC

[351] Nechemia Burshtein. On Natural Exactly Covering Systems of congruences having moduli occurring at most twice. Journal of Number Theory, 8(3):251–259, August 1976. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X76900019.

Warmus:1976:SNI

[352] M. Warmus. A supplementary note on the irregularities of distributions. Journal of Number Theory, 8(3):260–263, August 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X76900020.

Szemeredi:1976:PPE

[353] E. Szemerédi. On a problem of P. Erdős. Journal of Number Theory, 8 (3):264–270, August 1976. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X76900032.

Kisilevsky:1976:NFC

[354] H. Kisilevsky. Number fields with class number congruent to 4 mod 8 and Hilbert’s theorem 94. Journal of Number Theory, 8(3):271–279, August 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X76900044.

Chowla:1976:IPS

[355] P. Chowla and S. Chowla. On the integer points on some special hyperelliptic curves over a finite field. Journal of Number Theory, 8(3):280– 281, August 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X76900056.

Catlin:1976:TPMa

[356] Paul A. Catlin. Two problems in metric diophantine approximation, I. Journal of Number Theory, 8(3):282–288, August 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X76900068.

Catlin:1976:TPMb

[357] Paul A. Catlin. Two problems in metric diophantine approximation, II. Journal of Number Theory, 8(3):289–297, August 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X7690007X.

Divis:1976:LPS

[358] Bohuslav Divis. On the lattice points on strictly convex surfaces. Journal of Number Theory, 8(3):298–307, August 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X76900081.

Falk:1976:IIO

[359] Daniel Falk. On the invertibility of ideals in orders. Journal of Number Theory, 8(3):308–312, August 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X76900093.

Stephens:1976:PDSa

[360] P. J. Stephens. Prime divisors of second-order linear recurrences. I. Journal of Number Theory, 8(3):313–332, August 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X7690010X.

Stephens:1976:PDSb

[361] P. J. Stephens. Prime divisors of second order linear recurrences. II. Journal of Number Theory, 8(3):333–345, August 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X76900111.

Uchida:1976:CCF

[362] Kôji Uchida. On a cubic cyclic field with discriminant 163^{2}. Journal of Number Theory, 8(3):346–349, August 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X76900123.

Gruber:1976:BDM

[363] Peter Gruber. Eine Bemerkung über DOTU-Matrizen. (German) [A remark on DOTU-matrices]. Journal of Number Theory, 8(3):350–351, August 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X76900135.

Lang:1976:KZK

[364] H. Lang and R. Schertz. Kongruenzen zwischen Klassenzahlen quadratischer Zahlkörper. (German) [Congruences between class numbers of quadratic numbers]. Journal of Number Theory, 8(3):352–365, August 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X76900147.

Huxley:1976:NRP

[365] M. N. Huxley and K. Ramachandra. A note on recent papers of Ramachandra and Huxley. Journal of Number Theory, 8(3):366–368, August 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X76900159.

Anonymous:1976:PN

[366] Anonymous. Pages 369–496 (November 1976). Journal of Number Theory, 8(4):??, November 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Wunderlich:1976:PNT

[367] M. C. Wunderlich. The prime number theorem for random sequences. Journal of Number Theory, 8(4):369–371, November 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X76900846.

Richmond:1976:ARPa

[368] L. B. Richmond. Asymptotic results for partitions (I) and the distribution of certain integers. Journal of Number Theory, 8(4):372–389, November 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X76900858.

Richmond:1976:ARPb

[369] L. B. Richmond. Asymptotic results for partitions (II) and a conjecture of Bateman and Erdős. Journal of Number Theory, 8(4):390–396, November 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X7690086X.

Olson:1976:HIA

[370] Loren D. Olson. Hasse invariants and anomalous primes for elliptic curves with complex multiplication. Journal of Number Theory, 8(4):397–414, November 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X76900871.

Gold:1976:EIQ

[371] Robert Gold. Γ-extensions of imaginary quadratic fields. II. Journal of Number Theory, 8(4):415–419, November 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X76900883.

Gold:1976:IRI

[372] Robert Gold. Z_{3}-invariants of real and imaginary quadratic fields. Journal of Number Theory, 8(4):420–423, November 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X76900895.

Daily:1976:FIS

[373] Mary Lou Daily. A fundamental inequality for the sum of sets of m-tuples. Journal of Number Theory, 8(4):424–433, November 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X76900901.

Johnsen:1976:DIP

[374] John Johnsen. On the distribution of irregular primes. Journal of Number Theory, 8(4):434–437, November 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X76900913.

McQuillan:1976:GCD

[375] Donald L. McQuillan. On the Galois cohomology of Dedekind rings. Journal of Number Theory, 8(4):438–445, November 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X76900925.

Bernstein:1976:FUC

[376] Leon Bernstein. Fundamental units and cycles, I. Journal of Number Theory, 8(4):446–491, November 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X76900937.

Parry:1976:CB

[377] Charles J. Parry. On a conjecture of Brandler. Journal of Number Theory, 8(4):492–495, November 1976. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X76900949.

Anonymous:1976:AIV

[378] Anonymous. Author index for volume 8. Journal of Number Theory, 8 (4):496, November 1976. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X76900950.

Anonymous:1977:PF

[379] Anonymous. Pages 1–152 (February 1977). Journal of Number Theory, 9 (1):??, February 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1977:EB

[380] Anonymous. Editorial Board. Journal of Number Theory, 9(1):i, February 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X77900427.

Schwarz:1977:BDD

[381] W. Schwarz and H. Zassenhaus. Bohuslav Diviš: December 20, 1942–July 26, 1976. Journal of Number Theory, 9(1):iv–vii, February 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X77900439.

Moreno:1977:QPH

[382] Carlos J. Moreno. A quick proof of Hardy’s estimate for Ramanujan function τ(n). Journal of Number Theory, 9(1):1–3, February 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X77900440.

Furuya:1977:PIT

[383] Hisako Furuya. Principal ideal theorems in the genus field for absolutely abelian extensions. Journal of Number Theory, 9(1):4–15, February 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X77900452.

Liang:1977:MDS

[384] J. Liang and H. Zassenhaus. The minimum discriminant of sixth degree totally complex algebraic number fields. Journal of Number Theory, 9 (1):16–35, February 1977. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X77900464.

Gooding:1977:MFA

[385] Frederic Gooding. Modular forms arising from spherical polynomials and positive definite quadratic forms. Journal of Number Theory, 9(1):36–47, February 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X77900476.

Feldman:1977:ANA

[386] N. I. Feldman. Approximation of number π by algebraic numbers from special fields. Journal of Number Theory, 9(1):48–60, February 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X77900488.

Evans:1977:PSA

[387] Ronald Evans. Polynomial sums over automorphs of a positive definite binary quadratic form. Journal of Number Theory, 9(1):61–62, February 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X7790049X.

Olson:1977:ATF

[388] John E. Olson. An addition theorem for finite Abelian groups. Journal of Number Theory, 9(1):63–70, February 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X77900506.

Vaaler:1977:TTR

[389] Jeffrey D. Vaaler. A Tauberian theorem related to Weyl’s criterion. Journal of Number Theory, 9(1):71–78, February 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X77900518.

Weisman:1977:AD

[390] Carl S. Weisman. On p-adic differentiability. Journal of Number Theory, 9(1):79–86, February 1977. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X7790052X.

Loxton:1977:APC

[391] J. H. Loxton and A. J. van der Poorten. Arithmetic properties of certain functions in several variables. Journal of Number Theory, 9(1):87–106, February 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X77900531.

Newman:1977:TI

[392] Donald J. Newman. Tesselation of integers. Journal of Number Theory, 9 (1):107–111, February 1977. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X77900543.

Ozeki:1977:AFI

[393] Michio Ozeki. An application of L-functions of imaginary quadratic fields. Journal of Number Theory, 9(1):112–120, February 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X77900555.

Schecker:1977:MZM

[394] Hanno Schecker. Über die Menge der Zahlen, die als Minima quadratischer Formen auftreten. (German) [On the set of numbers that appear as minima of quadratic forms]. Journal of Number Theory, 9(1):121–141, February 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X77900567.

Cook:1977:DIM

[395] P. J. Cook. Diophantine inequalities with mixed powers. Journal of Number Theory, 9(1):142–152, February 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X77900579.

Anonymous:1977:PM

[396] Anonymous. Pages 153–278 (May 1977). Journal of Number Theory, 9 (2):??, May 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Mohanty:1977:SCD

[397] S. P. Mohanty. A system of cubic diophantine equations. Journal of Number Theory, 9(2):153–159, May 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X77900178.

Rosen:1977:ADC

[398] Michael Rosen. Ambiguous divisor classes in function fields. Journal of Number Theory, 9(2):160–174, May 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X7790018X.

Washington:1977:C

[399] Lawrence C. Washington. The calculation of L_{p}(1, X). Journal of Number Theory, 9(2):175–178, May 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X77900191.

Vegh:1977:PRC

[400] Emanuel Vegh. k-th-power residue chains. Journal of Number Theory, 9(2):179–181, May 1977. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X77900208.

Callahan:1977:FLK

[401] Thomas Callahan, Morris Newman, and Mark Sheingorn. Fields with large Kronecker constants. Journal of Number Theory, 9(2):182–186, May 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X7790021X.

Diaconis:1977:SOT

[402] Persi Diaconis, Frederick Mosteller, and Hironari Onishi. Second-order terms for the variances and covariances of the number of prime factors — Including the square free case. Journal of Number Theory, 9(2):187– 202, May 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X77900221.

Matthews:1977:EPR

[403] K. R. Matthews. An example from power residues of the critical problem of Crapo and Rota. Journal of Number Theory, 9(2):203–208, May 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X77900233.

Rosen:1977:SSD

[404] Kenneth H. Rosen. On the sign of some Dedekind sums. Journal of Number Theory, 9(2):209–212, May 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X77900245.

Shyr:1977:GDU

[405] Jih-Min Shyr. A generalization of Dirichlet’s unit theorem. Journal of Number Theory, 9(2):213–217, May 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X77900257.

Madden:1977:QFF

[406] Daniel J. Madden. Quadratic function fields with invariant class group. Journal of Number Theory, 9(2):218–228, May 1977. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X77900269.

Agou:1977:FCF

[407] Simon Agou. Factorisation sur un corps fini F_{p^{n}} des polynômes composés f(X^{p^{r}} −aX) lorsque f(X) est un polynôme irréductible de F_{p^{n}}(X). (French) [Factorization over a finite field F_{p^{n}} of the compound polynomials f(X^{p^{r}} −aX) when f(X) is a irreducible polynomial of F_{p^{n}}(X)]. Journal of Number Theory, 9(2):229–239, May 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X77900270.

Chakerian:1977:CBC

[408] G. D. Chakerian and H. Groemer. On convex bodies containing a given number of lattice points. Journal of Number Theory, 9(2):240–246, May 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X77900282.

Cohn:1977:RCG

[409] Harvey Cohn. Rational compositum genus for a pure cubic field. Journal of Number Theory, 9(2):247–257, May 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X77900294.

Robertson:1977:PLM

[410] M. M. Robertson. Partitions of large multipartites with congruence conditions. II. Journal of Number Theory, 9(2):258–270, May 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X77900300.

Turjanyi:1977:MAN

[411] S. Turjányi. On maximal asymptotic nonbases of zero density. Journal of Number Theory, 9(2):271–275, May 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X77900312.

Stephens:1977:CCC

[412] N. M. Stephens. On a conjecture of Chowla and Chowla. Journal of Number Theory, 9(2):276–277, May 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X77900324.

Parry:1977:CUA

[413] Charles J. Parry. Correction to “Units of Algebraic Numberfields”. Journal of Number Theory, 9(2):278, May 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X77900336. See [321].

Anonymous:1977:PA

[414] Anonymous. Pages 279–392 (August 1977). Journal of Number Theory, 9(3):??, August 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Jehne:1977:KCA

[415] Wolfram Jehne. Kronecker classes of algebraic number fields. Journal of Number Theory, 9(3):279–320, August 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X7790066X.

Grossman:1977:SRU

[416] Edward H. Grossman. Sums of roots of unity in cyclotomic fields. Journal of Number Theory, 9(3):321–329, August 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X77900671.

Bateman:1977:EFT

[417] Paul T. Bateman. The Erdős–Fuchs theorem on the square of a power series. Journal of Number Theory, 9(3):330–337, August 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X77900683.

Galambos:1977:NLP

[418] Janos Galambos. The number of large prime divisors of consecutive integers. Journal of Number Theory, 9(3):338–341, August 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X77900695.

Perlis:1977:E

[419] Robert Perlis. On the equation ζ_{k}(s) = ζ_{k′}(s). Journal of Number Theory, 9(3):342–360, August 1977. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X77900701.

Pinzur:1977:DDS

[420] Laurence Pinzur. Denominators of Dedekind sums. Journal of Number Theory, 9(3):361–369, August 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X77900713.

Schlickewei:1977:NFE

[421] Hans Peter Schlickewei. On norm form equations. Journal of Number Theory, 9(3):370–380, August 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X77900725.

Schlickewei:1977:LFA

[422] Hans Peter Schlickewei. On linear forms with algebraic coefficients and diophantine equations. Journal of Number Theory, 9(3):381–392, August 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X77900737.

Anonymous:1977:PN

[423] Anonymous. Pages 393–564 (November 1977). Journal of Number Theory, 9(4):??, November 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Jones:1977:QGQ

[424] Burton W. Jones. Quasi-genera of quadratic forms. Journal of Number Theory, 9(4):393–412, November 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X77900014.

Heidrich:1977:FDP

[425] Arnošt J. J. Heidrich. On finite differences, permutations, and Quadratic residues. Journal of Number Theory, 9(4):413–419, November 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X77900026.

Erdos:1977:BSI

[426] P. Erdős and D. J. Newman. Bases for sets of integers. Journal of Number Theory, 9(4):420–425, November 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X77900038.

Divis:1977:LPT

[427] Bohuslav Diviš. Lattice point theory in polyhedra. Journal of Number Theory, 9(4):426–435, November 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X7790004X.

Alladi:1977:DBP

[428] Krishnaswami Alladi. Duality between prime factors and an application to the prime number theorem for arithmetic progressions. Journal of Number Theory, 9(4):436–451, November 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X77900051.

Alladi:1977:DPP

[429] Krishnaswami Alladi and Charles Grinstead. On the decomposition of n! into prime powers. Journal of Number Theory, 9(4):452–458, November 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X77900063.

Pohst:1977:RTR

[430] Michael Pohst. Regulatorabschätzungen für total reelle algebraische Zahlkörper. (German) [Regulatory estimates for totally real algebraic numbers]. Journal of Number Theory, 9(4):459–492, November 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X77900075.

Srinivasan:1977:FLS

[431] S. Srinivasan. A footnote to the large sieve. Journal of Number Theory, 9 (4):493–498, November 1977. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X77900087.

Bremner:1977:SPS

[432] Andrew Bremner. Solution of a problem of Skolem. Journal of Number Theory, 9(4):499–501, November 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X77900099.

Chowla:1977:NZD

[433] S. Chowla, J. Cowles, and M. Cowles. On the number of zeros of diagonal cubic forms. Journal of Number Theory, 9(4):502–506, November 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X77900105.

Metsankyla:1977:SPN

[434] Tauno Metsänkylä. A short proof for the nonvanishing of a character sum. Journal of Number Theory, 9(4):507–509, November 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X77900117.

Hensley:1977:PWT

[435] Douglas Hensley. Polynomials which take Gaussian integer values at Gaussian integers. Journal of Number Theory, 9(4):510–524, November 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X77900129.

Richmond:1977:PEC

[436] L. B. Richmond. A problem of Erdős concerning partitions. Journal of Number Theory, 9(4):525–534, November 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X77900130.

Hinz:1977:NTA

[437] Jürgen Hinz. Über Nullstellen der m-ten Ableitung der Dedekindschen Zetafunktion. (German) [On zeroing the m-th derivative of the Dedekind zeta function]. Journal of Number Theory, 9(4):535–560, November 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X77900142.

Greaves:1977:CLP

[438] George Greaves. Correction to “Large prime factors of binary forms”. Journal of Number Theory, 9(4):561–562, November 1977. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:

//www.sciencedirect.com/science/article/pii/0022314X77900154.

See [63].

Anonymous:1977:AIV

[439] Anonymous. Author index for volume 9. Journal of Number Theory, 9 (4):563–564, November 1977. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X77900166.

Anonymous:1978:PF

[440] Anonymous. Pages 1–134 (February 1978). Journal of Number Theory, 10(1):??, February 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1978:EB

[441] Anonymous. Editorial Board. Journal of Number Theory, 10(1):i, February 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X7890001X.

Davis:1978:CNT

[442] Daniel Davis. Computing the number of totally positive circular units which are squares. Journal of Number Theory, 10(1):1–9, February 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X78900021.

vonLienen:1978:QF

[443] Horst von Lienen. The quadratic form x^{2} − 2py^{2}. Journal of Number Theory, 10(1):10–15, February 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X78900033.

Rudman:1978:GBU

[444] Ronald J. Rudman and Ray P. Steiner. A generalization of Berwick’s unit algorithm. Journal of Number Theory, 10(1):16–34, February 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X78900045.

Lundell:1978:DPS

[445] Albert T. Lundell. A divisibility property for Stirling numbers. Journal of Number Theory, 10(1):35–54, February 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X78900057.

Burde:1978:AMG

[446] Klaus Burde. Über Anzahlformeln modulo p. (German) [On number formulas modulo p]. Journal of Number Theory, 10(1):55–61, February 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X78900069.

Craig:1978:IV

[447] Maurice Craig. Integer values of ∑(x^{2}/yz). Journal of Number Theory, 10(1):62–63, February 1978. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X78900070.

Agou:1978:IPC

[448] Simon Agou. Irréductibilité des polynômes f(X^{p^{2r}} − aX^{p^{r}} − bX) sur un corps fini F_{p^{s}}. (French) [Irreducibility of polynomials f(X^{p^{2r}} −aX^{p^{r}} −bX) on a finite field F_{p^{s}}]. Journal of Number Theory, 10(1):64–69, February 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X78900082.

Schertz:1978:TRI

[449] Reinhard Schertz. Zur theorie der Ringklassenkörper über imaginärquadratischen Zahlkörpern. (German) [On the theory of ring class fields over imaginary-square number fields]. Journal of Number Theory, 10 (1):70–82, February 1978. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X78900094.

Graham:1978:AER

[450] S. Graham. An asymptotic estimate related to Selberg’s sieve. Journal of Number Theory, 10(1):83–94, February 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X78900100.

Erdos:1978:NIS

[451] P. Erdős and S. L. Segal. A note on Ingham’s summation method. Journal of Number Theory, 10(1):95–98, February 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X78900112.

Ratliff:1978:TSR

[452] Michael Ratliff. The Thue–Siegel–Roth–Schmidt theorem for algebraic functions. Journal of Number Theory, 10(1):99–126, February 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X78900124.

Cordes:1978:RSI

[453] Craig M. Cordes. Representation sets for integral binary quadratic forms over the rationals. Journal of Number Theory, 10(1):127–134, February 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X78900136.

Anonymous:1978:PM

[454] Anonymous. Pages 135–272 (May 1978). Journal of Number Theory, 10 (2):??, May 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Cohn:1978:FRC

[455] Harven Cohn. Formal ring of a cubic solid angle. Journal of Number Theory, 10(2):135–150, May 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X7890032X.

Schmidt:1978:DAF

[456] Asmus L. Schmidt. Diophantine approximation in the field Q(i(11^{1/2})). Journal of Number Theory, 10(2):151–176, May 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X78900331.

Grekos:1978:RDS

[457] Georges Grekos. Répartition des densités des sous-suites d’une suite d’entiers. Journal of Number Theory, 10(2):177–191, May 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X78900343.

Lew:1978:PIIa

[458] John S. Lew and Arnold L. Rosenberg. Polynomial indexing of integer lattice-points I. General concepts and quadratic polynomials. Journal of Number Theory, 10(2):192–214, May 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X78900355.

Lew:1978:PIIb

[459] John S. Lew and Arnold L. Rosenberg. Polynomial indexing of integer lattice-points II. Nonexistence results for higher-degree polynomials. Journal of Number Theory, 10(2):215–243, May 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X78900367.

Setzer:1978:UTC

[460] Bennett Setzer. Units in totally complex S_{3} fields. Journal of Number Theory, 10(2):244–249, May 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X78900379.

Hensley:1978:APS

[461] Douglas Hensley. An almost-prime sieve. Journal of Number Theory, 10(2):250–262, May 1978. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X78900380.

Odlyzko:1978:EPS

[462] Andrew M. Odlyzko and Richard P. Stanley. Enumeration of power sums modulo a prime. Journal of Number Theory, 10(2):263–272, May 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X78900392.

Anonymous:1978:PA

[463] Anonymous. Pages 273–383 (August 1978). Journal of Number Theory, 10(3):??, August 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Masley:1978:FFC

[464] John Myron Masley. On the first factor of the class number of prime cyclotomic fields. Journal of Number Theory, 10(3):273–290, August 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X78900240.

Coquet:1978:UDM

[465] Jean Coquet. On the uniform distribution modulo one of some subsequences of polynomial sequences. Journal of Number Theory, 10(3):291– 296, August 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X78900252.

Miki:1978:RBB

[466] Hiroo Miki. A relation between Bernoulli numbers. Journal of Number Theory, 10(3):297–302, August 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X78900264.

Madden:1978:AGA

[467] Daniel J. Madden. Arithmetic in generalized Artin–Schreier extensions of k(x). Journal of Number Theory, 10(3):303–323, August 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X78900276.

Odoni:1978:RAI

[468] R. W. K. Odoni. Representations of algebraic integers by binary quadratic forms and norm forms from full modules of extension fields. Journal of Number Theory, 10(3):324–333, August 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X78900288.

Morrison:1978:AAN

[469] John F. Morrison. Approximation of p-adic numbers by algebraic numbers of bounded degree. Journal of Number Theory, 10(3):334–350, August 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X7890029X.

Chowla:1978:PR

[470] P. Chowla and S. Chowla. On k-th power residues. Journal of Number Theory, 10(3):351–353, August 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X78900306.

Bernstein:1978:AU

[471] Leon Bernstein. Applications of units. Journal of Number Theory, 10 (3):354–383, August 1978. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X78900318.

Anonymous:1978:PN

[472] Anonymous. Pages 385–524 (November 1978). Journal of Number Theory, 10(4):??, November 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Newman:1978:FMD

[473] D. J. Newman and Ian Richards. Finite measures designed for accuracy on arithmetic progressions. Journal of Number Theory, 10(4):385–394, November 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X78900148.

Parson:1978:ABM

[474] Alayne Parson and Jack Tull. Asymptotic behavior of multiplicative functions. Journal of Number Theory, 10(4):395–420, November 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X7890015X.

Srinivasan:1978:TDH

[475] S. Srinivasan. On two-dimensional Hammersley’s sequences. Journal of Number Theory, 10(4):421–429, November 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X78900161.

Erdos:1978:DSI

[476] P. Erdős and A. Sárközy. On differences and sums of integers, I. Journal of Number Theory, 10(4):430–450, November 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X78900173.

Erdos:1978:CRP

[477] Paul Erdős, D. E. Penney, and Carl Pomerance. On a class of relatively prime sequences. Journal of Number Theory, 10(4):451–474, November 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X78900185.

Trinks:1978:BBV

[478] Wolfgang Trinks. Über B. Buchbergers Verfahren, Systeme algebraischer Gleichungen zu Lösen. (German) [On B. Buchberger’s method of solving systems of algebraic equations]. Journal of Number Theory, 10(4):475–488, November 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X78900197.

Perlis:1978:CNA

[479] Robert Perlis. On the class numbers of arithmetically equivalent fields. Journal of Number Theory, 10(4):489–509, November 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X78900203.

Metsankyla:1978:IIK

[480] Tauno Metsänkylä. Iwasawa invariants and Kummer conguruences. Journal of Number Theory, 10(4):510–522, November 1978. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X78900215.

Anonymous:1978:A

[481] Anonymous. Announcement. Journal of Number Theory, 10(4):523, November 1978. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X78900227.

Anonymous:1978:AIV

[482] Anonymous. Author index for volume 10. Journal of Number Theory, 10(4):524, November 1978. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X78900239.

Anonymous:1979:PF

[483] Anonymous. Pages 1–159 (February 1979). Journal of Number Theory, 11(1):??, February 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1979:EB

[484] Anonymous. Editorial Board. Journal of Number Theory, 11(1):i, February 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X79900131.

Jarden:1979:CLF

[485] Moshe Jarden and Jürgen Ritter. On the characterization of local fields by their absolute Galois groups. Journal of Number Theory, 11(1):1–13, February 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X79900143.

Pass:1979:PQA

[486] Ralph P. Pass. On the partial quotients of algebraic integers. Journal of Number Theory, 11(1):14–15, February 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X79900155.

Rickert:1979:SAG

[487] Udo Rickert. Über Summen in abelschen Gruppen. (German) [On sums in Abelian groups]. Journal of Number Theory, 11(1):16–19, February 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X79900167.

Agou:1979:IPC

[488] Simon Agou. Irréductibilité des polynômes f(X^{p^{2r}} − aX^{p^{r}} − bX) sur un corps fini F_{p^{s}}. (French) [Irreducibility of polynomials f(X^{p^{2r}} −aX^{p^{r}} −bX) on a finite field F_{p^{s}}]. Journal of Number Theory, 11(1):20, February 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X79900179.

Gillard:1979:RUC

[489] Roland Gillard. Remarques sur les unités cyclotomiques et les unités elliptiques. (French) [notes on cyclotomic units and elliptic units]. Journal of Number Theory, 11(1):21–48, February 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X79900180.

Cook:1979:DIM

[490] R. J. Cook. Diophantine inequalities with mixed powers, II. Journal of Number Theory, 11(1):49–68, February 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X79900192.

Erbach:1979:PGG

[491] D. W. Erbach, J. Fischer, and J. McKay. Polynomials with PSL(2, 7) as Galois group. Journal of Number Theory, 11(1):69–75, February 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X79900209.

Halberstam:1979:RRR

[492] H. Halberstam and H.-E. Richert. On a result of R. R. Hall. Journal of Number Theory, 11(1):76–89, February 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X79900210.

Ryavec:1979:NR

[493] C. Ryavec. A new representation of ζ(s). Journal of Number Theory, 11(1):90–94, February 1979. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X79900222.

Myerson:1979:NZD

[494] Gerald Myerson. On the number of zeros of diagonal cubic forms. Journal of Number Theory, 11(1):95–99, February 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X79900234.

Spackman:1979:SSD

[495] Kenneth W. Spackman. Simultaneous solutions to diagonal equations over finite fields. Journal of Number Theory, 11(1):100–115, February 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X79900246.

Langmann:1979:EFA

[496] Klaus Langmann. Eine endliche Formel für die Anzahl der Teiler von n. (German) [A finite formula for the number of divisors of n]. Journal of Number Theory, 11(1):116–127, February 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X79900258.

Schmidt:1979:GRA

[497] Claus-G. Schmidt. Grössencharaktere und Relativklassenzahl abelscher Zahlkörper. (German) [Size characters and relative class number of abelian number fields]. Journal of Number Theory, 11(1):128–159, February 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X7990026X.

Anonymous:1979:PM

[498] Anonymous. Pages 161–284 (May 1979). Journal of Number Theory, 11 (2):??, May 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Robert:1979:CE

[499] G. Robert. Caractères exceptionnels. Journal of Number Theory, 11 (2):161–170, May 1979. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X79900362.

Kimura:1979:KKV

[500] Noriaki Kimura. Kummersche Kongruenzen für die verallgemeinerten Bernoullischen Zahlen. (German) [Kummer’s congruences for generalized Bernoulli numbers]. Journal of Number Theory, 11(2):171–187, May 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X79900374.

Pfeuffer:1979:CAC

[501] Horst Pfeuffer. On a conjecture about class numbers of totally positive quadratic forms in totally real algebraic number fields. Journal of Number Theory, 11(2):188–196, May 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X79900386.

Erdos:1979:BNS

[502] Paul Erdős and Melvyn B. Nathanson. Bases and nonbases of square-free integers. Journal of Number Theory, 11(2):197–208, May 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X79900398.

Shallit:1979:SCF

[503] Jeffrey Shallit. Simple continued fractions for some irrational numbers. Journal of Number Theory, 11(2):209–217, May 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X79900404.

Arenstorf:1979:UDI

[504] R. F. Arenstorf and David Johnson. Uniform distribution of integral points on 3-dimensional spheres via modular forms. Journal of Number Theory, 11(2):218–238, May 1979. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X79900416.

Bloom:1979:ISE

[505] John R. Bloom. On the invariants of some Z_{i}-extensions. Journal of Number Theory, 11(2):239–256, May 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X79900428.

Erdos:1979:DOI

[506] P. Erdős and A. M. Odlyzko. On the density of odd integers of the form (p − 1)2^{−n} and related questions. Journal of Number Theory, 11(2):257– 263, May 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X7990043X.

Takacs:1979:GDS

[507] Lajos Takács. On generalized Dedekind sums. Journal of Number Theory, 11(2):264–272, May 1979. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X79900441.

Guting:1979:RAH

[508] Rainer Güting. Rationale Approximation mit Hilfe pythagoräischer Zahlen. (German) [Rational approximation using Pythagorean numbers]. Journal of Number Theory, 11(2):273–278, May 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X79900453.

Weinstein:1979:ZAS

[509] Lenard Weinstein. The zeros of the Artin l-series of a cubic field on the critical line. Journal of Number Theory, 11(2):279–284, May 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X79900465.

Anonymous:1979:PA

[510] Anonymous. Pages 285–475 (August 1979). Journal of Number Theory, 11(3):??, August 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Bambah:1979:D

[511] R. P. Bambah. Dedication. Journal of Number Theory, 11(3):285, August 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X79900015.

Ayoub:1979:SC

[512] Raymond Ayoub. Sarvadaman Chowla. Journal of Number Theory, 11 (3):286–301, August 1979. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X79900027.

Hockney:1979:RNP

[513] Donald Hockney. The relevance of natural philosophy for contemporary science. Journal of Number Theory, 11(3):302–304, August 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X79900039.

Goldfeld:1979:MVF

[514] D. Goldfeld and C. Viola. Mean values of L-funtions associated to elliptic, Fermat and other curves at the centre of the critical strip. Journal of Number Theory, 11(3):305–320, August 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X79900040.

Friedlander:1979:DRV

[515] John Friedlander and Stephen Halperin. Distinct representatives, varieties and rational homotopy. Journal of Number Theory, 11(3):321–323, August 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X79900052.

Smith:1979:DKS

[516] Robert A. Smith. On n-dimensional Kloosterman sums. Journal of Number Theory, 11(3):324–343, August 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X79900064.

Gross:1979:ICS

[517] Benedict H. Gross. On an identity of Chowla and Selberg. Journal of Number Theory, 11(3):344–348, August 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X79900076.

Berndt:1979:SGJ

[518] Bruce C. Berndt and Ronald J. Evans. Sums of Gauss, Jacobi, and Jacobsthal. Journal of Number Theory, 11(3):349–398, August 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X79900088.

Cohn:1979:QCG

[519] Harvey Cohn. Quaternionic compositum genus. Journal of Number Theory, 11(3):399–411, August 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X7990009X.

Lubkin:1979:FGL

[520] Saul Lubkin. Finite generation of lifted p-adic homology with compact supports. Generalization of the Weil conjectures to singular, noncomplete algebraic varieties. Journal of Number Theory, 11(3):412–464, August 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X79900106.

Ramachandra:1979:SRT

[521] K. Ramachandra. Some remarks on a theorem of Montgomery and Vaughan. Journal of Number Theory, 11(3):465–471, August 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X79900118.

Taussky:1979:DPA

[522] Olga Taussky. A diophantine problem arising out of similarity classes of integral matrices. Journal of Number Theory, 11(3):472–475, August 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X7990012X.

Anonymous:1979:PN

[523] Anonymous. Pages 477–560 (November 1979). Journal of Number Theory, 11(4):??, November 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Newman:1979:WPR

[524] D. J. Newman and Morton Slater. Waring’s problem for the ring of polynomials. Journal of Number Theory, 11(4):477–487, November 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X79900271.

Bhaskaran:1979:CGF

[525] M. Bhaskaran. Construction of genus field and some applications. Journal of Number Theory, 11(4):488–497, November 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:

//www.sciencedirect.com/science/article/pii/0022314X79900283.

See comment [533].

Gordon:1979:A

[526] Basil Gordon and Murray Schacher. The admissibility of A_{5}. Journal of Number Theory, 11(4):498–504, November 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X79900295.

Cook:1979:NCS

[527] R. J. Cook. A note on character sums. Journal of Number Theory, 11 (4):505–515, November 1979. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X79900301.

Cook:1979:SPI

[528] Roger Cook. On sums of powers of integers. Journal of Number Theory, 11 (4):516–528, November 1979. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X79900313.

Gerstein:1979:UQF

[529] Larry J. Gerstein. Unimodular quadratic forms over global function fields. Journal of Number Theory, 11(4):529–541, November 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X79900325.

Elliott:1979:TID

[530] P. D. T. A. Elliott and P. Erdős. The tails of infinitely divisible laws and a problem in number theory. Journal of Number Theory, 11(4):542–551, November 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X79900337.

Brunotte:1979:ETK

[531] Horst Brunotte and Franz Halter-Koch. Zur Einheitenberechnung in totalreellen kubischen Zahlkörpern nach Godwin. (German) [On unit calculation in totally real cubic numbers according to Godwin]. Journal of Number Theory, 11(4):552–559, November 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X79900349.

Anonymous:1979:AIV

[532] Anonymous. Author index for volume 11. Journal of Number Theory, 11(4):560, November 1979. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X79900350.

Anonymous:1984:CGF

[533] Anonymous. “Construction of genus field and some applications” by M. Bhaskaran. Journal of Number Theory, 19(3):449–451, December 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84900842. See [525].